### v2

# MaskCLIP++: A Mask-Based CLIP Fine-tuning Framework for Open-Vocabulary Image Segmentation

Abstract
Open-vocabulary image segmentation has been advanced through the synergy between mask generators and vision-language models like Contrastive Language-Image Pre-training (CLIP). Previous approaches focus on generating masks while aligning mask features with text embeddings during training. In this paper, we observe that relying on generated low-quality masks can weaken the alignment of vision and language in regional representations. This motivates us to present a new fine-tuning framework, named MaskCLIP++, which uses ground-truth masks instead of generated masks to enhance the mask classification capability of CLIP. Due to the limited diversity of image segmentation datasets with mask annotations, we propose incorporating a consistency alignment constraint during fine-tuning, which alleviates categorical bias toward the fine-tuning dataset. After low-cost fine-tuning, combining with the mask generator in previous state-of-the-art mask-based open vocabulary segmentation methods, we achieve performance improvements of +1.7, +2.3, +2.1, +3.1, and +0.3 mIoU on the A-847, PC-459, A-150, PC-59, and PAS-20 datasets, respectively. Code is available at https://github.com/HVision-NKU/MaskCLIPpp.
1 Introduction
Image segmentation is one of the most extensively studied tasks in computer vision, which aims to partition an image into several regions where each region corresponds to the same object or shares consistent semantics. Traditional image segmentation models are often defined on a closed vocabulary. When new classes need to be segmented, it typically requires redefining the vocabulary set, annotating data, and retraining the model. However, the cost of annotating image segmentation on a large vocabulary is high. Therefore, the task of Open-Vocabulary Segmentation (OVS) has been proposed to perform image segmentation under any vocabulary set.
Large-scale image-text pre-training models, such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), have demonstrated strong zero-shot recognition capabilities, sparking interest in transferring these abilities to image segmentation tasks. Several studies (Ding et al., 2023; Jiao et al., 2023; Yu et al., 2024; Jiao et al., 2024) have explored how to adapt mask generators (Cheng et al., 2021; 2022) and pre-trained CLIP to each other to achieve open-vocabulary segmentation at both the semantic and instance levels.
However, we argue that continuing to invest in improving the mask generation may not be cost-effective. For different CLIP architectures, even when the generated masks closely match the ground truth (GT) masks, the improvement in segmentation performance is limited to no more than 5% as shown in Figure 1(a). Conversely, for different mask generators, if CLIP can classify the generated masks as accurately as the GT class assignments, the segmentation performance improvement can exceed 40% as shown in Figure 1(b). Furthermore, Figure 1(a) shows that for the original CLIP, there remains a performance gap of 3.0%–4.7% between the generated masks and the GT masks, implying that continuing to use the generated masks during training may negatively impact the alignment quality between regional images and text. We confirm this inference in our experiments.
The above analysis indicates that GT masks can replace mask generators during fine-tuning CLIP with image segmentation data, which not only eliminates unnecessary training costs but also improves the alignment quality between regional images and text. However, when attempting to use image segmentation data to fine-tune CLIP, a significant challenge is to avoid overfitting to the training set. Existing approaches either integrate a frozen CLIP during inference (Xu et al., 2022b; 2023a; Yu et al., 2024) or employ a frozen teacher model for distillation (Jiao et al., 2023; 2024). The former may limit the model’s performance improvement, while the latter often requires careful tuning of distillation strategies and hyperparameters. To address this, we analyze the sources of misalignment in previous methods and propose a consistency alignment constraint to ensure that alignment optimization preserves CLIP’s original embedding space. The outcome of our explorations is a new CLIP fine-tuning framework, coined as MaskCLIP++.
Experiments show that MaskCLIP++ can significantly improve the mask classification capabilities of pretrained CLIP across multiple architectures, achieving 10.1%-16.8% mIoU improvement on ADE20K (Zhou et al., 2017). Additionally, MaskCLIP++ can be used with different mask generators to achieve open-vocabulary segmentation at the semantic- or instance-level. Compared to recent state-of-the-art methods, MaskCLIP++ offers lower training cost in terms of training time and memory usage, while achieving even better performance. We hope that the universality and efficiency of MaskCLIP++ could inspire further research in open-vocabulary image segmentation.
(a) Improvement potential of mask generation
(b) Improvement potential of mask classification
Figure 1: (a) Performance comparison of various original CLIP architectures using the same generated and ground truth masks. (b) Performance comparison of various mask generators with different backbones using CLIP based on ConvNeXt-L and ground truth classification. All models are trained on the COCO Panoptic dataset (Lin et al., 2014), with results reported on the ADE-150 (Zhou et al., 2017) dataset, measured by mean Intersection over Union (mIoU).
2 Related Work
Mask-based segmentation. Mask-based segmentation is a technique that views image segmentation as the process of mask generation and mask classification, which is also known as region-based segmentation. To address instance and panoptic segmentation, He et al. (2017); Wang et al. (2020; 2021); Zhang et al. (2021) proposed various mask generators. Recently, Cheng et al. (2021) unified the semantic, panoptic and instance segmentation into a single architecture. MaskFormer (Cheng et al., 2021) uses learnable queries to interact with image features, which are then decoded as masks and class scores. Mask2Former (Cheng et al., 2022) further accelerates the convergence of MaskFormer and reduces memory usage through multi-scale decoder design, using masks from shallow decoders as priors for deeper layers, and employing sampling-based optimization. Current mask-based OVS methods need almost to retrain an open-vocabulary mask generator based on Mask2Former (Xu et al., 2023d; a; Yu et al., 2024). In contrast, our method can directly use the Mask2Former trained on a closed vocabulary during inference, reducing training cost and increasing flexibility in usage.
Open-vocabulary segmentation. Depending on the data used, existing methods have developed into different paradigms. The first paradigm offers many heuristic solutions based on observations of CLIP without additional training (Zhou et al., 2022; Wang et al., 2024a; Li et al., 2023; Sun et al., 2024; Lan et al., 2024a; b). The second paradigm utilizes image-text pair data (Thomee et al., 2016; Changpinyo et al., 2021) to train OVS models under weak supervision (Xu et al., 2022a; Luo et al., 2023; Xu et al., 2023b; Zhang et al., 2023; Mukhoti et al., 2023; Cha et al., 2023; Liu et al., 2024a). Due to the lack of precise segmentation locations, these methods suffer from poor segmentation quality. The third paradigm enhances segmentation quality by incorporating image-mask data (Kirillov et al., 2023) and knowledge from SAM (Wang et al., 2024c; Yuan et al., 2024). However, the use of SAM not only increases inference cost but also relies on users or other detectors to provide high-quality prompts. The fourth paradigm seeks to achieve mutual adaptation between the segmentation model and CLIP through fully supervised training on image segmentation datasets containing image, mask, and category triplets, thereby producing an open-vocabulary segmentation model. Based on the type of the segmentation model, these methods can be categorized into mask-based approaches (Ding et al., 2022; Xu et al., 2022b; Ghiasi et al., 2022; Han et al., 2023; Liang et al., 2023; Qin et al., 2023; Jiao et al., 2023; Ding et al., 2023; Xu et al., 2023d; c; a; Yu et al., 2024; Liu et al., 2024b; Jiao et al., 2024) and pixel/patch-based ones (Li et al., 2022; Zhou et al., 2023; Li et al., 2024; Cho et al., 2024; Xie et al., 2024). Pixel-based OVS methods typically focus on semantic segmentation and cannot perform instance-level segmentation independently. Additionally, some unique approaches to achieving OVS involve image retrieval (Shin et al., 2022; Barsellotti et al., 2024), feature distillation (Chen et al., 2023; Wu et al., 2024) or model merging (Wang et al., 2024b). Some unified large models also incorporate the functionality of OVS (Zou et al., 2023; Shen et al., 2024). This paper focuses on efficiently transferring CLIP on limited segmentation datasets to improve its performance on semantic- and instance-level open-vocabulary segmentation tasks.
3 Methodology
3.1 Revisiting Image-Text Embeddings Alignment for OVS
We first revisit the roles of the visual encoder (CLIP-V) and text encoder (CLIP-T) in the pretrained CLIP model. Denote the image embeddings from CLIP-V as $E_{i}\in\mathbb{R}^{1\times D}$, where $D$ represents the dimension of embedding space. To create text embeddings, a vocabulary set of $K$ categories is transformed into the corresponding sentences using templates like “A photo of {}”, and then encoded by CLIP-T, yielding $E_{t}\in\mathbb{R}^{K\times D}$. The original CLIP model maps both images and text into an aligned embedding space. In this space, image-text pairs maintain semantic alignment, which we describe as “pre-alignment” and can often be measured using cosine similarity $\langle E_{i},E_{t}\rangle$. In open-vocabulary segmentation, preserving the pre-alignment property is crucial for adapting CLIP to zero-shot dense predictions, as it enhances the alignment of region-specific image and text embeddings.
Previous mask-based OVS methods obtain pre-aligned mask embeddings and text embeddings in two ways. As illustrated in Figure 2(a), the first way aims to train a mask generator that adapts to CLIP, with the generalization mainly derived from the original frozen parameters of CLIP (Xu et al., 2023c; Ding et al., 2023). However, attempts to align the embedding spaces of the mask generator and CLIP often fail (Xu et al., 2022b; Yu et al., 2024). As illustrated in Figure 2(b), the second way focuses on making CLIP sensitive to the generated masks during training to enhance performance and avoids overfitting through distillation (Jiao et al., 2023; 2024).
We identify two limitations of these two types of methods. First, they rely on noisy generated masks, which can degrade the quality of the region representations (as shown by the vertical comparisons in Figure 1(a)). This degradation can weaken the alignment between region-specific image and text embeddings. Second, there is a lack of an intuitive explanation and a straightforward solution for learning better alignment without severe overfitting. Moreover, freezing CLIP limits its transferability from global to local features, while distillation typically requires carefully designed strategies and incurs additional training cost.
To overcome these limitations, we propose a new training pipeline for OVS, named MaskCLIP++, which, to our knowledge, is the first method that attempts to leverage ground truth (GT) masks to obtain region features instead of relying on the mask generator during training. Furthermore, we propose the constraint named “consistency alignment” to alleviate overfitting when adding extra parameters to optimize alignment. We will give detailed descriptions in what follows.
(a) Adapt Mask Generator to CLIP
(b) Adapt CLIP to Mask Generator
(c) MaskCLIP++
Figure 2: Comparison of training pipeline between previous mask-based OVS methods. $\mathcal{L}$ denotes the loss function. (2(a)) adapts the mask generator to CLIP and avoids overfitting by freezing CLIP. (2(b)) adapts CLIP to the mask generator and avoids overfitting by distillation. “T” and “S” denote the teacher and student models, respectively. Our method (2(c)) abandons the mask generator and avoids overfitting by consistency alignment.
3.2 A New Fine-tuning Framework for CLIP
In this subsection, we detail how to use our framework to fine-tune CLIP and integrate it with an off-the-shelf mask generator. The overall framework is illustrated in Figure 3.
Extracting mask embeddings. Disregarding architectural details, CLIP-V can be considered as consisting of two parts: feature extraction by the encoder and feature aggregation by the neck. The mask is inserted after the last feature $F_{n}\in\mathbb{R}^{C\times H\times W}$ of the encoder is obtained, converting the original global aggregation into an aggregation within the mask’s effective region as follows:
$E_{m}=\operatorname{Proj}\sum_{i,j}\left({\phi(M)}_{i,j}\cdot{(F_{n})}_{:,i,j}\right),$ (1)
where $M$ is an object mask that has been downsampled to the same size as $F_{n}$, $E_{m}\in\mathbb{R}^{1\times D}$ is the embedding representations of this mask, $i,j$ are the spatial coordinates, and $\operatorname{Proj}$ represents the formulation of the neck that implements channel projection. The weighting function $\phi$ is related to the original CLIP pooling structure. Specifically, if the original CLIP uses global average pooling, $\phi$ acts as an L1 normalization function to achieve mask average pooling. If the original CLIP uses attention pooling, $\phi$ can be expressed as follows:
$\phi(M)=\operatorname{Softmax}\left(\frac{q(T_{\text{cls}})k(T_{ij})}{\sqrt{d}}+\alpha M_{\text{bin}}\right),$ (2)
where $T_{\text{cls}}$ and $T_{ij}$ represent the global and local representations sent to the attention pooling, respectively. $q,k$ are the input projection of the attention. $d$ is the hidden size of attention. The spatially expanded and binarized version of $M$ is denoted as $M_{\text{bin}}$, where any elements less than $\max(M)/2$ are set to $\infty$- ∞.
Parameterized similarity modeling with consistency alignment. The pretrained CLIP mask embedding $E_{m}$ and the category text embedding $E_{t}$ are initially pre-aligned with the initial similarity map $S=\langle E_{m},E_{t}\rangle$ being relatively coarse. Optimizing this alignment with some supervision is needed to achieve a more accurate similarity map. We refer to this process as parameterized similarity modeling (PSM).
However, when incorporating the mask generator into similarity modeling, the mask embeddings produced by the mask generator align almost exclusively with CLIP’s text embeddings on the training set, resulting in the misalignment illustrated in Figure 2(a). We first analyze the cause of this overfitting to develop our similarity modeling approach.
To facilitate discussion, we simplify this approach as follows. Let $t\in\mathbb{R}^{D}$ denote a normalized text embedding from CLIP, and let $m_{1},m_{2}\in\mathbb{R}^{D}$ denote different mask embeddings. The initial similarity is defined as $s=m^{T}t$. Now, before computing the inner product across modalities, the mask embeddings are updated using a parameter $\Theta\in\mathbb{R}^{D\times D}$, resulting in a new similarity $r=(\Theta m)^{T}t$. Although $r$ can be optimized to be close to the target, an unintuitive phenomenon may arise: After mask embeddings are updated via gradient descent, the order of $r_{1}$ and $r_{2}$ is closer to the target, but the order of $s_{1}$ and $s_{2}$ may deviate further from the target. This phenomenon is illustrated by a toy example in Appendix Figure 5.
Consequently, the original similarity relationships between CLIP’s embeddings are disrupted. Our experiments demonstrate that the additional parameter $\Theta$ fails to achieve CLIP’s level of generalization when fine-tuning on limited image segmentation data, resulting in overfitting. Therefore, when attempting to parameterize similarity modeling, it is essential to ensure that the modeled similarity maintains the same order as the original similarity, so that both can be optimized toward the same target. We refer to this constraint as “consistency alignment”.
To optimize under the consistency alignment constraint in most of the time, the simplest way to achieve this is to update the similarity map in a dimension that is irrelevant to any modality using linear layers. Specifically, for $Q$ masks and $K$ texts, the $Q\times K$ similarity matrix $S$ is first expanded to $Q\times K\times T$ using a linear layer of size $T$, and then it is reduced to $Q\times K$ using another linear layer of the same size.
When using the image patch embeddings $E_{p}$ to construct pseudo text embeddings (Jiao et al., 2024), a residual connection to the original text embeddings is necessary to ensure consistency alignment to some extent, as shown below:
$S=\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle,\quad\text{where}~{}P_{t}=\operatorname{Attention}(q(E_{t}),k(E_{p}),v(E_{p}))$ (3)
Here, $\operatorname{Attention}$ denotes the scaled dot-product attention (Vaswani, 2017), $q,k,v$ denote the input projection functions for attention and $\operatorname{Norm}$ denotes the layer norm (Lei Ba et al., 2016). The image patch embeddings $E_{p}$ are equivalent to specialized mask embeddings extracted from unit masks on a per-patch basis.
Figure 3: The detailed framework of MaskCLIP++. The PSM represents the parameterized similarity modeling, which is constrained by the consistency alignment. The mask generator is only used during inference, and can be flexibly replaced.
Integrating with off-the-shelf mask generators. As the training process is entirely decoupled from the mask generator, we can flexibly employ various types of mask generators during inference to achieve semantic- or instance-level open-vocabulary segmentation. Moreover, integration with the class probabilities of the mask generator can appropriately improve the mask classification performance in the training categories. Specifically, let $\mathcal{K}$ be the training vocabularies of the mask generator. $P_{c}$ and $P_{s}$ are the class probabilities of the fine-tuned CLIP and the mask generator, respectively. The integrated class probabilities $P_{\gamma}$ can be expressed as follows:
$P_{\gamma}^{(i)}=\begin{cases}{P_{s}^{(i)}}^{\gamma}\cdot{P_{c}^{(i)}}^{(1-\gamma)},&\text{if}~{}i\in\mathcal{K}\\ P_{c}^{(i)}.&\text{otherwise}\\ \end{cases}$ (4)
Compared to previous OVS methods, our MaskCLIP++ offers the following several advantages. First, we use GT masks during training, which are less noisy and more accurate, thus improving alignment quality and reducing training cost. Second, from the perspective of avoiding overfitting, the consistency alignment constraint provides more optimization space than freezing CLIP (Xu et al., 2023c; Yu et al., 2024), and does not rely on carefully designed distillation strategies (Wu et al., 2024; Jiao et al., 2024), resulting in faster convergence with a single optimization objective.
4 Experiments
4.1 Implementation Details
In the following experiments, we utilize various architectures of pre-trained CLIP to demonstrate the generality of our method. The CLIP models with ResNet architecture are sourced from OpenAI (Radford et al., 2021), those with ViT architecture from EVA-CLIP (Sun et al., 2023), and those with ConvNeXt architecture from OpenCLIP (Cherti et al., 2023).
By default, consistent with previous work on universal open-vocabulary semantic segmentation (Yu et al., 2024), we fine-tune our model using the COCO Panoptic training set (Lin et al., 2014), which consists of 118K images and 133 categories. The AdamW optimizer is utilized with an initial learning rate of 2e-4, decaying to 1e-5 via cosine annealing. The fine-tuning process is conducted for 20K iterations with a batch size of 4.
The size of the training image is adjusted according to the downsampling factor of the model to ensure that the final size of the feature map is $32^{2}$. The parameters involved in fine-tuning differ slightly across architectures. For the ConvNeXt and ResNet architectures, all parameters of CLIP-V are fine-tuned. For the ViT architecture, following Cho et al. (2024), only the positional encodings and embedding layers of CLIP-V are fine-tuned. In the ConvNeXt architecture, the fine-tuning learning rate for CLIP parameters is 1e-3 times that of other parts, while in the ResNet and ViT architectures, it is 1e-2 times.
During inference, we mainly evaluate using the ADE20K validation set (Zhou et al., 2017), which contains 2,000 images and 150 categories, along with annotations for semantic segmentation, panoptic segmentation, and instance segmentation. In ADE20K, 64 categories overlap with those in COCO, while the remaining 86 categories do not. We measure semantic segmentation performance using Intersection over Union (IoU), recording mIoU(S) for overlapping categories, mIoU(U) for non-overlapping categories, and overall mIoU. Panoptic Quality (PQ) and Average Precision (AP) are used to assess performance in panoptic segmentation and instance segmentation, respectively.
By default, we use FC-CLIP (Yu et al., 2024) trained on the COCO Panoptic (Lin et al., 2014) training set as the mask generator. For the base and large level CLIP, we employ the corresponding FC-CLIP models as mask generators  [^1] Since FC-CLIP did not provide a mask generator with CLIP ConvNeXt-B as the backbone, we used their code for CLIP ConvNeXt-L to produce a base one.. The inference size of the mask generator remains consistent with FC-CLIP, set to 800. The inference size of CLIP depends on the model architecture, and the input sizes and initial segmentation performance for different architectures can be seen in Table 4.
[…]
PSM C.A. mIoU(S) mIoU(U) mIoU
Baseline - 32.7 16.3 23.3
$\langle\operatorname{Linear}(E_{m}),E_{t}\rangle$ 32.4 1.9 14.9
$\langle E_{m},\operatorname{Linear}(E_{t})\rangle$ 32.5 1.8 14.9
$\operatorname{Linear}\langle E_{m},E_{t}\rangle$ 45.4 25.7 34.1
$\langle E_{m},\operatorname{Norm}(P_{t})\rangle$ 31.9 4.5 16.2
$\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle$ 46.3 26.7 35.1
Table 2: Impact of prior quality during training. “Seg masks” is the matched generated masks. “GT boxes” is the minimum bounding rectangles of GT masks.
Prior Type mIoU(S) mIoU(U) mIoU
Pixels 44.3 24.2 32.8
Seg masks 44.9 23.7 32.7
GT boxes 45.4 25.8 34.2
GT masks 46.3 26.7 35.1
4.2 Ablation Study of MaskCLIP++
In our ablation study, we used CLIP ConvNeXt-B along with the mask generator from FC-CLIP (Yu et al., 2024), which shares the same backbone. The direct combination of them achieved a baseline mIoU of 23.3 on ADE20K (Zhou et al., 2017).
Importance of consistency alignment. We use different parameterized similarity modeling methods to illustrate the importance of the consistent alignment constraint. As shown in Table 2, the attempts to use linear parameters to update mask embeddings or text embeddings violate the constraint of consistency alignment during fine-tuning, resulting in a significant performance drop on unseen categories compared to the baseline (1.9% vs. 16.3% or 1.8% vs. 16.3%). The $P_{t}$ are weighted sums of image patch embeddings as shown in equation 3, so $\langle E_{m},\operatorname{Norm}(P_{t})\rangle$ cannot have a pre-aligned initial similarity map, leading to substantial performance drops on unseen categories as well (4.5% vs. 16.3%). The PSM of $\operatorname{Linear}\langle E_{m},E_{t}\rangle$ adheres to the consistency alignment constraint, resulting in performance improvements on both seen and unseen categories (45.4% vs. 32.7% and 25.7% vs. 16.3%). For $\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle$, the image context information $P_{t}$ indirectly interacts with $E_{m}$. This compensates for the rigidity of mask average pooling in the CLIP ConvNeXt architecture, ultimately resulting in improved performance (35.1% vs. 34.1%). When fine-tuning CLIP models with attention pooling (ViT and ResNet), we still use the PSM of $\operatorname{Linear}\langle E_{m},E_{t}\rangle$. We present additional experiments in Appendix Table 12 to demonstrate the extent to evaluate how the PSMs in Table 2 preserve or disrupt the consistency alignment constraint.
Impact of prior quality during training. We refer to the regional information fed into CLIP during the mask embedding extraction stage as “priors” and investigate how the quality of these priors affects fine-tuning performance. We experiment with pixels, generated masks that match GT masks according to IoU, GT masks, and the GT boxes as their minimum bounding rectangles. As shown in Table 2, using GT masks can achieve the best performance. Using the generated masks as prior yields lower performance compared to the GT masks (32.7% vs. 35.1%), indicating that incorporating the mask generator during training actually hinders CLIP fine-tuning. Both GT masks and GT boxes priors outperform pixel priors, suggesting that fine-tuning CLIP from global recognition to region-level recognition is less challenging than fine-tuning it to pixel-level recognition.
Table 3: Impact of different mask generators. “OV” indicates that the mask generator utilizes a pre-trained CLIP as the backbone and employs CLIP text embeddings as a variable-length classification head, typically considered an open-vocabulary mask generator. $P_{s}$, $P_{c}$ and $P_{\gamma}$ are defined in equation 4. mIoU are reported. Latency (ms) is the average inference time of a single image on ADE20K, measured in an environment with Pytorch 2.3.1, cuda 11.8 and one RTX 2080Ti GPU.
Mask Generator + MaskCLIP++
Backbone Decoder OV $P_{s}$ $P_{c}$ $P_{\gamma=0.1}$ $P_{\gamma=0.2}$ $P_{\gamma=0.4}$ Latency (ms)
Swin-T Mask2Former 15.7 33.9 34.3 34.2 33.5 15.5
Swin-L Mask2Former 16.8 33.9 34.3 33.9 33.2 33.6
ConvNeXt-B FC-CLIP 19.1 35.1 35.7 35.8 35.0 32.3
ConvNeXt-L FC-CLIP 20.8 34.6 35.4 35.8 35.8 39.0
Table 4: Improvements of various CLIP architectures. All models are fine-tuned on COCO Panoptic and evaluated on ADE20K. “Res.” denotes the input resolution of CLIP during inference. For CNN-based CLIP, all parameters in CLIP-V are fine-tuned. For ViT-based CLIP, only the position embedding and input projections of the attention in the CLIP-V are fine-tuned. For RN50x16, ConvNeXt-B, and ViT-B/16, use the mask generator with ConvNeXt-B backbone. For others, use the mask generator with ConvNeXt-L backbone. Ensemble strategy is not used.
CLIP Fine-tuned Input Original CLIP MaskCLIP++
Arch. Part Params(M) Res mIoU PQ AP mIoU PQ AP
RN50x16 V-All 167.3 512 18.4 15.8 9.4 29.3+10.9 21.8+6.0 11.1+1.7
ConvNeXt-B V-All 90.3 512 23.3 19.1 11.6 35.1+11.8 24.5+5.4 13.6+2.0
ConvNeXt-L V-All 202.7 640 25.4 20.7 14.6 35.6+10.2 26.5+5.8 16.7+2.1
ConvNeXt-XXL V-All 846.5 640 26.3 20.3 13.3 36.4+10.1 27.1+6.8 16.6+3.3
ViT-B/16 V-Attn 14.3 384 17.0 15.0 9.5 33.8+16.8 24.4+9.4 13.2+3.7
ViT-L/14 V-Attn 50.9 392 20.3 17.4 12.4 36.6+16.3 27.3+9.9 17.0+4.6
ViT-G/14 V-Attn 238.3 336 20.7 18.6 13.5 36.8+16.1 27.7+9.1 17.1+3.6
Impact of different mask generators. Various types of mask generators can be integrated with our fine-tuned CLIP, as shown in Table 3. We use mask generators from Mask2Former (Cheng et al., 2022) and FC-CLIP (Yu et al., 2024), which are typically considered as closed-vocabulary and open-vocabulary mask generators, respectively. These mask generators are all trained on COCO Panoptic (Lin et al., 2014) and evaluated on ADE20K (Zhou et al., 2017) using mIoU. As shown in Table 3, open-vocabulary mask generators do not significantly outperform closed-vocabulary mask generators in terms of segmentation performance ($P_{s}$). And using our fine-tuned CLIP for mask classification can significantly improve the segmentation performance ($P_{c}$) of all mask generators. This suggests that both open-vocabulary and closed-vocabulary mask generators possess an inherent category-independent mask generation capability that has yet to be fully harnessed. Furthermore, integrating the classification scores of the mask generator on training categories, as described in equation 4, can slightly enhance performance. For the hyperparameter $\gamma$, empirically, the worse the mask classification quality $P_{s}$ of the mask generator is, the smaller $\gamma$ can be set.
4.3 Enhancement of Mask Classification by MaskCLIP++
We apply our method to CLIP with various architectures to demonstrate its universality. The experimental settings are described in Section 4.1. Table 4 presents the types of CLIP architectures, the parts that were fine-tuned, and the corresponding number of fine-tuned parameters. Compared to the original CLIP, our method improved mIoU by 10.1%-16.8%, PQ by 5.4%-9.9% and AP by 1.7%-4.6% on ADE20K(Zhou et al., 2017) when using COCO Panoptic (Lin et al., 2014) for fine-tuning. Additionally, we demonstrate significant performance improvements of the original CLIP across different input resolutions, as shown in Figure 6 in Appendix C.
Table 5: Comparisons with previous methods on Open-Vocabulary Semantic Segmentation task.
Method VLM Training dataset A-847 PC-459 A-150 PC-59 PAS-20
Non-mask-based OVS method
SED (Xie et al., 2024) CLIP ConvNeXt-B COCO-Stuff 11.4 18.6 31.8 57.7 94.4
CAT-Seg (Cho et al., 2024) CLIP ViT-B/16 COCO-Stuff 12.0 19.0 31.8 57.5 94.6
Mask-based OVS method
ZSSeg (Xu et al., 2022b) CLIP ViT-B/16 COCO-Stuff 7.0 - 20.5 47.7 -
OpenSeg (Ghiasi et al., 2022) ALIGN EN-B7 COCO-Captions + Loc. Narr. 8.8 12.2 28.6 48.2 72.2
DeOP (Han et al., 2023) CLIP ViT-B/16 COCO-Stuff-156 7.1 9.4 22.9 48.8 91.7
FreeSeg (Qin et al., 2023) CLIP ViT-B/16 COCO-Panoptic - - 24.6 - -
MAFT (Jiao et al., 2023) CLIP ViT-B/16 COCO-Stuff 10.1 12.8 29.1 53.5 90.0
SAN (Xu et al., 2023c) CLIP ViT-B/16 COCO-Stuff 10.1 12.6 27.5 53.8 94.0
SCAN (Liu et al., 2024b) CLIP ViT-B/16 COCO-Stuff 10.8 13.2 30.8 58.4 97.0
FC-CLIP (Yu et al., 2024) CLIP ConvNeXt-B COCO-Panoptic 13.1 16.5 31.2 55.2 94.2
MAFT+ (Jiao et al., 2024) CLIP ConvNeXt-B COCO-Stuff 13.8 16.2 34.6 57.5 95.4
MaskCLIP++ CLIP ConvNeXt-B COCO-Stuff 14.5 18.7 35.4 59.1 95.8
Non-mask-based OVS method
OVSeg (Liang et al., 2023) CLIP ViT-L/14 COCO-Stuff 9.0 12.4 29.6 55.7 94.5
CLIPSelf (Wu et al., 2024) CLIP ViT-L/14 COCO-Stuff 12.4 - 34.5 62.3 -
SED (Xie et al., 2024) CLIP ConvNeXt-L COCO-Stuff 13.7 22.1 35.2 60.6 96.1
CAT-Seg (Cho et al., 2024) CLIP ViT-L/14 COCO-Stuff 16.0 23.8 37.9 63.3 97.0
Mask-based OVS method
MaskCLIP (Ding et al., 2023) CLIP ViT-L/14 COCO-Panoptic 8.2 10.0 23.7 45.9 -
MasQCLIP (Xu et al., 2023d) CLIP ViT-L/14 COCO-Panoptic 10.7 18.2 30.4 57.8 -
ODISE (Xu et al., 2023a) CLIP ViT-L/14 COCO-Panoptic 11.1 14.5 29.9 57.3 -
SAN (Xu et al., 2023c) CLIP ViT-L/14 COCO-Stuff 12.4 15.7 32.1 57.7 94.6
SCAN (Liu et al., 2024b) CLIP ViT-L/14 COCO-Stuff 14.0 16.7 33.5 59.3 97.2
FC-CLIP (Yu et al., 2024) CLIP ConvNeXt-L COCO-Panoptic 14.8 18.2 34.1 58.4 95.4
MAFT+ (Jiao et al., 2024) CLIP ConvNeXt-L COCO-Stuff 15.1 21.6 36.1 59.4 96.5
MaskCLIP++ CLIP ViT-L/14 COCO-Stuff 16.8 23.9 38.2 62.5 96.8
4.4 Comparison with previous methods
Open-vocabulary semantic segmentation. There are notable differences in the training settings compared to Section 4.1 to be consistent with most previous open vocabulary segmentation works. We fine-tune CLIP on COCO-Stuff (Caesar et al., 2018), which shares the same images as COCO Panoptic (Lin et al., 2014) but defines 171 categories. We report mIoU performance on the validation sets of the following datasets: ADE-20K (Zhou et al., 2017) with 847 classes (A-847) and 150 classes (A-150), PASCAL-Context (Mottaghi et al., 2014) with 459 classes (PC-459) and 59 classes (PC-59), and PASCAL-VOC (Everingham et al., 2010) with 20 foreground classes (PAS-20).
For comparisons with previous works using base-level VLM models, we fine-tune CLIP ConvNeXt-B with previous settings. For comparisons with works using larger-level VLM models, we fine-tune CLIP ViT-L/14. Apart from the previous settings, we also fine-tune the input projections of the attention in the CLIP text encoder following CAT-Seg (Cho et al., 2024). We use the same mask generator as MAFT+ (Jiao et al., 2024) during inference and set $\gamma=0.4$ for CLIP ConvNeXt-B and $\gamma=0.1$ for CLIP ViT-L/14.
Table 5 shows that our base model achieves the best performance on A847, A150 and PC-59, and our large model achieves the best performance on A-847, PC-459, and A-150.
Table 6: Comparisons with previous methods on Open-Vocabulary Panoptic and Instance Segmentation tasks. All methods are trained on COCO Panoptic (Lin et al., 2014) and evaluated on ADE20K (Zhou et al., 2017).
Method VLM PQ AP
FreeSeg (Qin et al., 2023) CLIP ViT-B/16 16.3 6.5
FC-CLIP (Yu et al., 2024) CLIP ConvNeXt-B 23.1 13.1
MaskCLIP++ CLIP ConvNeXt-B 24.5 13.6
MaskCLIP (Ding et al., 2023) CLIP ViT-L/14 15.1 6.2
MasQCLIP (Xu et al., 2023d) CLIP ViT-L/14 23.3 -
ODISE (Xu et al., 2023a) CLIP ViT-L/14 22.6 14.4
FC-CLIP (Yu et al., 2024) CLIP ConvNeXt-L 26.8 16.8
MAFT+ (Jiao et al., 2024) CLIP ConvNeXt-L 27.1 -
MaskCLIP++ CLIP ViT-L/14 27.3 17.0
Table 7: Comparison of CLIP’s mask classification ability in different methods. The same ViT-L/14 architecture of CLIP and the same mask generator are used. The first three rows are initialized by OpenAI-CLIP, and the last three rows are initialized by EVA-CLIP.
Method PC-459 A-150 PC-59 Stuff
OpenAI 9.3 17.6 27.4 18.0
CAT-Seg 19.6 34.1 57.3 45.6
MaskCLIP++ 21.3 34.8 59.3 47.2
EVA02 11.1 20.1 26.0 18.3
CLIPSelf 13.0 22.9 39.1 27.1
MaskCLIP++ 24.2 37.8 62.3 50.6
Open-vocabulary panoptic and instance segmentation. Following the settings of previous works, we fine-tune CLIP on COCO Panoptic (Lin et al., 2014). All methods are evaluated on ADE20K, with PQ measuring panoptic segmentation performance and AP measuring instance segmentation performance. We use the same mask generator as FC-CLIP (Yu et al., 2024) and do not use the ensemble strategy ($\gamma=0.0$). Table 6 shows that our method achieves the best performance.
Inference with the same masks. Considering that some competitive methods in Table 5 do not explicitly use a mask generator and fine-tune CLIP in alternative ways, we test the mask classification performance of CLIPs from various sources under the same generated masks. Since the PSM module we trained may not be compatible with their CLIP, we remove the PSM during inference. The ensemble strategy with the mask generator is also excluded. As shown in Table 7, our fine-tuned CLIPs outperform ones from other methods in mask classification.
Inferece without mask generators. Because the goal of our fine-tuning is to enhance the mask classification ability of CLIP, which indirectly enhances the local representation between different patch representations of the CLIP. As shown in Table 8, after applying to some methods that do not require training and additional mask generators, further performance improvements can be obtained.
Table 8: Performance improvements achieved when combined with several unsupervised open-vocabulary semantic segmentation methods that do not rely on a mask generator. OpenAI CLIP ViT-B/16 are used. “Citys” means cityscapes dataset (Cordts et al., 2016).
Method A-847 PC-459 A-150 PC-59 Stuff Citys
SCLIP (Wang et al., 2024a) - 7.7 16.1 34.2 22.4 32.2
SCLIP w/ MaskCLIP++ - 9.5 19.3 40.8 29.0 35.9
ClearCLIP (Lan et al., 2024a) - 7.9 16.7 36.0 23.9 30.0
ClearCLIP w/ MaskCLIP++ - 9.6 19.2 40.8 30.4 33.5
ProxyCLIP  (Lan et al., 2024b) 7.4 8.4 20.2 39.1 26.5 38.1
ProxyCLIP w/ MaskCLIP++ 8.0 9.9 23.2 44.6 34.2 42.4
Trade-off between training cost and performance. Figure 4 shows the trade-off between training cost and performance of different open-vocabulary segmentation methods. As our method does not use the mask generator or teacher model during training but only a small number of GT masks instead of a large number of generated masks and has only one classification loss, it converges quickly, has lower memory usage, and achieves better performance.
Data efficiency. Since the goal of our fine-tuning is to transfer CLIP’s open-vocabulary recognition capabilities to mask regions, rather than learning category-specific pixel groupings from segmentation datasets, our method can effectively fine-tune using a smaller amount of image segmentation data. We randomly sampled varying proportions of data from the COCO-Stuff training set for fine-tuning, using the same configuration as the base model in Table 5. As shown in Table 9, even with only 0.1% of the data (approximately 118 images with annotations), the fine-tuned model still demonstrates competitive performance.
Table 9: The impact of the amount of data used for fine-tuning. “data size” indicates the proportion of data used for fine-tuning relative to the original COCO-Stuff training set, and the fine-tuned model is CLIP-ConvNeXt-B.
Data size A-847 PC-459 A-150 PC-59 Stuff
100% 14.5 18.7 35.4 59.1 47.1
10% 14.2 19.0 34.9 59.2 47.2
5% 13.9 18.7 35.2 59.6 47.3
1% 13.8 19.1 35.1 58.5 47.0
0.1% 12.8 18.6 34.1 58.2 46.1
Figure 4: Trade-off between training cost and performance of different OVS methods. Radius size represents the memory usage during training. Using CLIP with base size for comparison.
Qualitative comparison of segmentation results. Figure 7 shows the qualitative results on the Open-Vocabulary Semantic Segmentation task. Using the mask generator from MAFT+ (Jiao et al., 2024), we compare the segmentation results of the original CLIP, MAFT+, and our method. Compared to the original CLIP, our method achieves more accurate recognition of local images. Compared to MAFT+, our method retains more of the original CLIP’s knowledge without generating too many erroneous biases. Our method also predicts relatively correct semantics for some objects that are not annotated in the GT.
5 Conclusions
In this paper, we point that using generated masks hinders CLIP’s regional image-text alignment in previous OVS work. And there is limited room for improving mask generation compared to mask classification. We also reveal that attempts to update CLIP’s embedding vectors disrupt its original pre-alignment properties, resulting in new alignments that function only on the training data. Based on these analysis, we propose the MaskCLIP++, which utilizes ground truth masks as prior to fine-tune CLIP, and models similarity under a consistency alignment constraint to mitigate overfitting. Our method significantly enhances CLIP’s mask recognition capabilities across various architectures. By integrating mask generators from prior work, we achieve comparable or superior performance on semantic segmentation, panoptic segmentation, and instance segmentation tasks with lower training costs. Overall, we adapt CLIP more effectively to open-vocabulary segmentation tasks through the use of image segmentation data and lightweight fine-tuning.
Appendix
Appendix A Limitations and Future work
First, the image masks must be downsampled to match the size of CLIP’s final feature map, which undergoes a downsampling factor of 14 to 32 times. Consequently, some generated small object masks may not be recognizable by MaskCLIP++.
Second, the mask generators employed during inference incur negligible computational overhead. As shown in Table 3, heavier backbones seem to enhance classification capabilities of their own ($P_{s}$), but do not significantly improve open-vocabulary segmentation performance when combined with our fine-tuned CLIP ($P_{c}$). Therefore, lightweight, category-agnostic mask generators continue to represent a promising research direction.
Last, our reliance on GT masks somewhat limits our ability to scale up at the data level. Finding ways to utilize more image-text pairs or image-mask pairs for fine-tuning remains an open question.
Appendix B Datasets of training and evaluation
The image segmentation datasets we used to train are COCO Panoptic (Lin et al., 2014) or COCO-Stuff (Caesar et al., 2018). For evaluation, we utilize the ADE20K (Zhou et al., 2017), PASCAL VOC (Everingham et al., 2010), and PASCAL Context (Mottaghi et al., 2014) datasets. A brief introduction to these datasets is provided below.
COCO Panoptic. It contains 164K images with 133 annotated classes, comprising 80 “thing” categories and 53 “stuff” categories. The “thing” categories are annotated at the instance level, with different instances of the same category using separate masks within a single image. In contrast, the “stuff” categories are semantically annotated, where all regions belonging to the same category share a single mask in an image.
COCO Stuff. It shares the same training images with COCO Panoptic. These images are annotated for semantic segmentation into 171 categories.
ADE20K. It contains 2,000 images for validation with two types of annotations: one for panoptic or semantic segmentation with 150 categories (dubbed as A-150), and the other for semantic segmentation with 847 categories (dubbed as A-847). Among the 150 categories defined in ADE20K, 64 appear in COCO Panoptic, while 86 do not. The names of these categories are detailed in Table 10.
PASCAL VOC. It includes 1,449 images for testing with 20 annotated foreground classes for semnantic segmentation.
PASCAL Context. It is a dataset for semantic understanding that contains 5,000 validation images. Two versions are employed for semantic segmentation: one with 59 frequently used classes (referred to as PC-59) and another with the complete set of 459 classes (referred to as PC-459).
Table 10: Categories in ADE20K that are present in COCO Panoptic (seen) or are absent from COCO Panoptic (unseen).
Split Name
Seen (64) wall, building, sky, floor, tree, ceiling, road, bed, windowpane, grass, cabinet, sidewalk,
person, door, table, mountain, plant, curtain, chair, car, water, sofa, shelf, sea,
mirror, rug, fence, rock, lamp, counter, sand, sink, refrigerator, stairs, pillow, river,
bridge, toilet, flower, book, bench, palm tree, boat, bus, towel, light bulb, truck,
television receiver, airplane, apparel, bottle, tent, oven, food, microwave, plant pots,
animal, bicycle, blanket, vase, traffic light, plate, cup, clock
Unseen (86) earth, painting, house exterior, field, armchair, seat, desk, wardrobe,
bathtub, railing, cushion, pedestal, box, column, signboard, chest of drawers, skyscraper,
fireplace, grandstand, path, runway, case, pool table, screen door, stairway, bookcase,
window screen, coffee table, hill, countertop, stove, kitchen island, computer, swivel chair,
arcade machine, hovel, tower, chandelier, awning, streetlight, booth, dirt track, pole, land,
bannister, escalator, ottoman, buffet, poster, stage, van, ship, fountain, conveyer belt,
washer, plaything, swimming pool, stool, barrel, basket, waterfall, bag, minibike, cradle,
step, tank, trade name, lake, dishwasher, projection screen, sculpture, exhaust hood, sconce,
tray, ashcan, ceiling fan, pier, crt screen, monitor, bulletin board, shower, radiator,
canopy, flag, bar, ball
Appendix C Additional Experiments
C.1 Additional Illustrations
Implementation details of Figure 1 We use mask generation and mask classification as the inference modes, with Figure 1(a) and Figure 1(b) presenting the performance upper bounds of the ideal mask generator and ideal mask classifier, respectively. The ideal mask generator is implemented by directly using the ground truth masks. To implement the ideal mask classifier, we compute the maximum matching between the generated masks and ground truth masks based on IoU, then assign the corresponding GT categories to the matched masks and combine them into the output segmentation maps.
Figure 5: A toy example: Because the inconsistency alignment between $s$ (similarity before $\Theta$) and $r$ (similarity after $\Theta$), after a round of gradient descent, $s$ is further from the target $\hat{s}$.
Phenomenon of inconsistency alignment. We first use a toy example to illustrate the potential consequences of destroying alignment consistency. Assume that we use the parameter $\Theta$ to update the mask embeddings, changing the original similarity from $s$ to $r$, and then optimize towards the target. However, due to the lack of explicit constraints on $\Theta$, there may be a situation as shown in Figure 5, where the order of $s$ and $r$ is inconsistent, resulting in correct optimization on $r$ but incorrect optimization on $s$. For simplicity, $\mathcal{L}$ in Figure 5 denotes an L1 loss without softmax, and only mask embeddings $m$ are updated in one round of gradient descent.
Futuremore, we use Table 12 as a supplement to Table 2 to illustrate the extent to which different PSM structures destroy the consistency alignment constraint. For models fine-tuned with various PSMs, we test by retaining or removing the PSM module, and compare the performance differences to assess whether the alignment consistency is maintained before and after the PSM intervention. The results show that only the three blue PSM structures largely satisfy the consistency alignment constraint, while the three red PSM structures severely disrupt it.
Table 11: The phenomenon of consistency alignment.
PSM test mIoU(S) mIoU(U) mIoU
- - 32.7 16.3 23.3
$\langle\operatorname{Linear}(E_{m}),E_{t}\rangle$ 37.7 20.2 27.7
32.4 1.9 14.9
$\langle E_{m},\operatorname{Linear}(E_{t})\rangle$ 37.6 20.3 27.6
32.5 1.8 14.9
$\operatorname{Linear}\langle E_{m},E_{t}\rangle$ 45.2 24.6 33.4
45.4 25.7 34.1
$\langle E_{m},\operatorname{Norm}(P_{t})\rangle$ 39.5 21.4 29.1
31.9 4.5 16.2
$\langle E_{m},\operatorname{Norm}(E_{t})\rangle$ 45.8 24.4 33.5
46.2 25.2 34.1
$\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle$ 44.0 24.1 32.6
46.3 26.7 35.1
Table 12: Performance of MaskCLIP++ on closed-vocabulary image segmentation tasks.
CLIP arch. Dataset mIoU PQ AP
ConvNeXt-B COCO-Stuff 47.1 - -
ViT-L/14 COCO-Stuff 51.0 - -
ConvNeXt-B COCO-Panoptic 57.6 46.1 34.4
ViT-L/14 COCO-Panoptic 63.3 52.3 42.3
The performance of closed-vocabulary image segmentation. We present the performance of MaskCLIP++ on closed-vocabulary image segmentation in Table 12, with the same configuration as in Table 5 and Table 6, respectively.
C.2 Additional Ablations
(a) ConvNeXt-B
(b) ConvNeXt-L
(c) ViT-B/16
(d) ViT-L/14
Figure 6: Under various input sizes of different CLIP models, our method (blue bars) has significant improvements over the original CLIP (red bars). The input size of the mask generator remains unchanged. All models are fine-tuned on COCO Panoptic and evaluated on ADE20K with mIoU reported.
Impact of CLIP input size. Considering that input image size often has a significant impact on the mIoU of semantic segmentation models, we compared the performance of our method with the original CLIP under different input sizes. As shown in Figure 6, with a fixed input size for the mask generator, our fine-tuned CLIP consistently outperforms the original CLIP across various input sizes for open-vocabulary semantic segmentation.
Overall, segmentation performance initially increases and then decreases as input size grows. Since our fine-tuned CLIP can better recognize local image features, the optimal input size tends to be smaller than that of the original CLIP. Our framework allows us to maintain the input size for the mask generator while reducing the input size for mask recognition, thereby enhancing performance and reducing computational costs.
Impact of integrating image context information into mask representation. Averaging CLIP features over perfectly precise mask regions is not conducive to extracting mask representations with stronger recognition capabilities, as recognition tasks typically require contextual information and focused attention on target internal regions. Here, we distinguish how we address this common issue using CLIP’s original pooling method. For CLIP models utilizing Attention Pooling (e.g., ViT-B/16), the hyperparameter $\alpha$ controls the degree of contextual information incorporated in the mask representation. As shown in Table 1313(b), smaller $\alpha$ values, indicating more extensive use of contextual information, lead to better performance. For CLIP models employing global average pooling (e.g. ConvNeXt-B), as shown in Table 1313(a), using global image embeddings $E_{i}$ or patch embeddings $E_{p}$ to construct pseudo-text embeddings $P_{t}$ in the PSM stage enhances the utilization of image context information to some extent, thereby improving performance.
Choice of Loss. Table 15 presents the performance differences observed when fine-tuning CLIP ConvNeXt-B with various loss functions. Training with COCO Panoptic (Lin et al., 2014) and evaluating on ADE20K (Zhou et al., 2017). When the training labels for a single image include multiple instances of the same category, such as pixels or instance masks, we re-weight the loss by categories to prevent bias toward more common classes during optimization. We adopt a class-reweighted cross-entropy loss for all our experiments in the main text.
Position of the mask prior insertion. Due to the repetitive nature of the ViT architecture, the position of mask prior insertion in CLIP models based on such architectures is not absolute. Following the approach of Xu et al. (2023c) for constructing attention masks, we can insert the mask prior at an earlier position than the last attention block. Table 15 takes ViT-B/16 as an example, demonstrating that inserting the mask prior at the third-to-last layer yields better performance. Therefore, in Table 4, the mask prior is inserted at the third-to-last layer for all ViT-based CLIP models.
Table 13: Impact of integrating image context information into mask embeddings $E_{m}$. The implementation varies based on the original CLIP’s pooling method.
(a) The case of CLIP using global average pooling, taking ConvNeXt-B as an example. $\Psi$ is a placeholder for image context information, and $P_{t}$ defined in equation 3 can be rewritten as $P_{t}=\operatorname{Attn}(q(E_{t}),k(\Psi),v(\Psi))$. $P_{t}$ interacts with $E_{m}$ in the PSM stage to increase the utilization of image context information.
Method in PSM Image Context mIoU
No $\Phi$ 34.1
$\Phi=E_{i}$ 34.5
$\Phi=E_{p}$ 35.1
(b) The case of CLIP using attention pooling, taking ViT-B/16 as an example. $\alpha$ is defined in equation 2. The bigger $\alpha$ is, the less the semantic similarity of the image context is utilized. The learnable $\alpha$ is initialized from $e^{-5}$ and learned in the logarithmic space.
Method in CLIP Image Context mIoU
$\alpha=e^{5}$ 31.1
$\alpha=e^{-5}$ 33.5
learnable $\alpha$ 33.8
Table 14: Choice of Loss. “Re.” means the loss is reweighted by categories. CLIP ConvNeXt-B is used.
Prior Loss Re. mIoU(S) mIoU(U) mIoU
Pixel CE 39.5 21.2 29.0
CE 44.3 23.4 32.8
Instance Mask BCE 45.7 24.5 33.5
BCE 46.3 25.3 34.3
CE 45.5 26.0 34.4
CE 46.3 26.7 35.1
Table 15: The impact of the position of mask prior insertion for ViT-B/16 CLIP. “Prior pos.” denotes the insert position relative to the total number of attention blocks.
Prior pos. mIoU PQ
-1 33.3 24.1
-2 33.5 24.5
-3 33.8 24.4
-4 33.5 24.4
Qualitative results. Figure 7 compares the visual results of open-vocabulary semantic segmentation using different methods with the same mask generator. We show more qualitative results of semantic- and instance level of open-vocabulary segmentation in Figure 8.
Image                    CLIP                MAFT+              MaskCLIP++ GT
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: Visualizations of open-vocabulary semantic segmentation on ADE20K (Zhou et al., 2017). From left to right are the input image, the result using the original CLIP, the result of MAFT+ (Jiao et al., 2024), the result using MaskCLIP++, and the ground truth. The same mask generator is used for CLIP, MAFT+ and MaskCLIP++.
Image                Semantic Segmentation      Panoptic Segmentation Instance Segmentation
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 8: More visualizations of semantic- and instance-level of open-vocabulary segmentation. We use mask generator from FC-CLIP (Yu et al., 2024) with ConvNeXt-L backbone. And use CLIP ViT-L/14 fine-tuned on COCO Panoptic (Lin et al., 2014). Images are from ADE20K (Zhou et al., 2017). The confidience threshold for instance segmentation is 0.3.

### v3

@@
-MaskCLIP++: A Mask-Based CLIP Fine-tuning Framework for Open-Vocabulary Image Segmentation
+High-Quality Mask Tuning Matters for Open-Vocabulary Segmentation
 Abstract
@@
 This motivates us to present a new fine-tuning framework, named MaskCLIP++, which uses ground-truth masks instead of generated masks to enhance the mask classification capability of CLIP.
-Due to the limited diversity of image segmentation datasets with mask annotations, we propose incorporating a consistency alignment constraint during fine-tuning, which alleviates categorical bias toward the fine-tuning dataset.
-After low-cost fine-tuning, combining with the mask generator in previous state-of-the-art mask-based open vocabulary segmentation methods, we achieve performance improvements of +1.7, +2.3, +2.1, +3.1, and +0.3 mIoU on the A-847, PC-459, A-150, PC-59, and PAS-20 datasets, respectively.
-Code is available at https://github.com/HVision-NKU/MaskCLIPpp.
+Due to the limited diversity of image segmentation datasets with mask annotations, we propose incorporating a consistency alignment principle during fine-tuning, which alleviates categorical bias toward the fine-tuning dataset.
+After low-cost fine-tuning, MaskCLIP++ significantly improves the mask classification performance on multi-domain datasets.
+Combining with the mask generator in previous state-of-the-art mask-based open vocabulary segmentation methods, we achieve performance improvements of +1.7, +2.3, +2.1, +3.1, and +0.3 mIoU on the A-847, PC-459, A-150, PC-59, and PAS-20 datasets, respectively.
+Code is avaliable at https://github.com/HVision-NKU/MaskCLIPpp.
 1 Introduction
+(a) Performance: generated masks v.s. ground truth masks.
+(b) Performance: CLIP classification v.s. ground truth classification.
+Figure 1: Observations: 1(a) demonstrates the potential negative impact of low-quality generated masks on CLIP’s mask classification learning, while 1(b) showcases the untapped generalization potential of existing mask generators.
+Mask generators are trained on the COCO [36], with results reported on the ADE20K [80].
 Image segmentation is one of the most extensively studied tasks in computer vision, which aims to partition an image into several regions where each region corresponds to the same object or shares consistent semantics.
@@
 However, the cost of annotating image segmentation on a large vocabulary is high.
-Therefore, the task of Open-Vocabulary Segmentation (OVS) has been proposed to perform image segmentation under any vocabulary set.
-Large-scale image-text pre-training models, such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), have demonstrated strong zero-shot recognition capabilities, sparking interest in transferring these abilities to image segmentation tasks.
-Several studies (Ding et al., 2023; Jiao et al., 2023; Yu et al., 2024; Jiao et al., 2024) have explored how to adapt mask generators (Cheng et al., 2021; 2022) and pre-trained CLIP to each other to achieve open-vocabulary segmentation at both the semantic and instance levels.
-However, we argue that continuing to invest in improving the mask generation may not be cost-effective.
-For different CLIP architectures, even when the generated masks closely match the ground truth (GT) masks, the improvement in segmentation performance is limited to no more than 5% as shown in Figure 1(a).
-Conversely, for different mask generators, if CLIP can classify the generated masks as accurately as the GT class assignments, the segmentation performance improvement can exceed 40% as shown in Figure 1(b).
-Furthermore, Figure 1(a) shows that for the original CLIP, there remains a performance gap of 3.0%–4.7% between the generated masks and the GT masks, implying that continuing to use the generated masks during training may negatively impact the alignment quality between regional images and text.
-We confirm this inference in our experiments.
-The above analysis indicates that GT masks can replace mask generators during fine-tuning CLIP with image segmentation data, which not only eliminates unnecessary training costs but also improves the alignment quality between regional images and text.
-However, when attempting to use image segmentation data to fine-tune CLIP, a significant challenge is to avoid overfitting to the training set.
-Existing approaches either integrate a frozen CLIP during inference (Xu et al., 2022b; 2023a; Yu et al., 2024) or employ a frozen teacher model for distillation (Jiao et al., 2023; 2024).
+Large-scale image-text pre-training models, such as CLIP [47] and ALIGN [25], have demonstrated strong zero-shot recognition capabilities, sparking interest in transferring these abilities to image segmentation tasks.
+Several studies [15, 26, 76, 27] have explored how to adapt mask generators [9, 10] and pre-trained CLIP to each other to achieve open-vocabulary segmentation at both the semantic and instance levels.
+However, our analysis based on two key observations reveals critical shortcomings in existing research that have been consistently overlooked.
+First, we establish a baseline method enabling CLIP to perceive mask categories by referencing prior methodologies [73, 76].
+To evaluate the impact of mask quality on inference performance, we conduct comparative experiments using generated versus ground-truth masks across various pre-trained CLIP models.
+As shown in Fig.
+1(a), generated masks exhibit inferior performance compared to ground-truth masks due to their low segmentation accuracy, semantic ambiguity, and excessive redundancy.
+Our experiments further demonstrate that employing higher-quality masks during training consistently yields superior performance under identical inference protocols.
+Second, to validate the strong generalization capability of existing mask generators, we evaluate multiple architectures pre-trained on the COCO dataset [36] and evaluate on the ADE20K dataset [80] with a different vocabulary set.
+Through IoU-based bipartite matching, we reassign “ground-truth categories” to predicted masks to determine whether these generators produce masks for unseen categories.
+As illustrated in Fig.
+1(b), the performance could be substantially improved if mask classification were perfect.
+This discrepancy suggests that existing mask generators inherently produce numerous valid masks for unseen categories, but their potential is severely limited by erroneous mask classification.
+Based on these observations, we propose to eliminate the reliance on mask generators during training and directly fine-tune CLIP using high-quality image segmentation data.
+However, since image segmentation data is typically defined on closed vocabularies, mitigating the issue of CLIP overfitting the training data after fine-tuning poses an important challenge.
+Existing approaches either integrate a frozen CLIP during inference [72, 71, 76] or employ a frozen teacher model for distillation [26, 27].
 The former may limit the model’s performance improvement, while the latter often requires careful tuning of distillation strategies and hyperparameters.
-To address this, we analyze the sources of misalignment in previous methods and propose a consistency alignment constraint to ensure that alignment optimization preserves CLIP’s original embedding space.
+To address this issue more effectively, we analyze the causes of modality misalignment in prior works and propose that the introduction of additional parameters should adhere to the consistent alignment principle, thereby preserving CLIP’s original embedding space throughout the alignment optimization process.
 The outcome of our explorations is a new CLIP fine-tuning framework, coined as MaskCLIP++.
-Experiments show that MaskCLIP++ can significantly improve the mask classification capabilities of pretrained CLIP across multiple architectures, achieving 10.1%-16.8% mIoU improvement on ADE20K (Zhou et al., 2017).
+Experimental results demonstrate that MaskCLIP++ can significantly enhance CLIP’s masked classification capability across datasets spanning diverse domains.
 Additionally, MaskCLIP++ can be used with different mask generators to achieve open-vocabulary segmentation at the semantic- or instance-level.
 Compared to recent state-of-the-art methods, MaskCLIP++ offers lower training cost in terms of training time and memory usage, while achieving even better performance.
-We hope that the universality and efficiency of MaskCLIP++ could inspire further research in open-vocabulary image segmentation.
-(a) Improvement potential of mask generation
-(b) Improvement potential of mask classification
-Figure 1: (a) Performance comparison of various original CLIP architectures using the same generated and ground truth masks.
-(b) Performance comparison of various mask generators with different backbones using CLIP based on ConvNeXt-L and ground truth classification.
-All models are trained on the COCO Panoptic dataset (Lin et al., 2014), with results reported on the ADE-150 (Zhou et al., 2017) dataset, measured by mean Intersection over Union (mIoU).
+We hope that the universality and efficiency of MaskCLIP++ could inspire further research in open-vocabulary segmentation.
 2 Related Work
@@
 Mask-based segmentation is a technique that views image segmentation as the process of mask generation and mask classification, which is also known as region-based segmentation.
-To address instance and panoptic segmentation, He et al.
-(2017); Wang et al.
-(2020; 2021); Zhang et al.
-(2021) proposed various mask generators.
-Recently, Cheng et al.
-(2021) unified the semantic, panoptic and instance segmentation into a single architecture.
-MaskFormer (Cheng et al., 2021) uses learnable queries to interact with image features, which are then decoded as masks and class scores.
-Mask2Former (Cheng et al., 2022) further accelerates the convergence of MaskFormer and reduces memory usage through multi-scale decoder design, using masks from shallow decoders as priors for deeper layers, and employing sampling-based optimization.
-Current mask-based OVS methods need almost to retrain an open-vocabulary mask generator based on Mask2Former (Xu et al., 2023d; a; Yu et al., 2024).
+To address instance and panoptic segmentation, [22, 62, 60, 79] proposed various mask generators.
+Recently, [9] unified the semantic, panoptic and instance segmentation into a single architecture.
+MaskFormer [9] uses learnable queries to interact with image features, which are then decoded as masks and class scores.
+Mask2Former [10] further accelerates the convergence of MaskFormer and reduces memory usage through multi-scale decoder design, using masks from shallow decoders as priors for deeper layers, and employing sampling-based optimization.
+Current mask-based OVS methods need almost to retrain an open-vocabulary mask generator based on Mask2Former [74, 71, 76].
 In contrast, our method can directly use the Mask2Former trained on a closed vocabulary during inference, reducing training cost and increasing flexibility in usage.
+(a) Adapt Mask Generator to CLIP
+(b) Adapt CLIP to Mask Generator
+(c) MaskCLIP++
+Figure 2: Comparison of training pipeline between previous mask-based OVS methods.
+$\mathcal{L}$ denotes the loss function.
+2(a) adapts the mask generator to CLIP and avoids overfitting by freezing CLIP.
+2(b) adapts CLIP to the mask generator and avoids overfitting by distillation.
+“T” and “S” denote the teacher and student models, respectively.
+Our method 2(c) abandons the mask generator and avoids overfitting by consistency alignment.
 Open-vocabulary segmentation.
 Depending on the data used, existing methods have developed into different paradigms.
-The first paradigm offers many heuristic solutions based on observations of CLIP without additional training (Zhou et al., 2022; Wang et al., 2024a; Li et al., 2023; Sun et al., 2024; Lan et al., 2024a; b).
-The second paradigm utilizes image-text pair data (Thomee et al., 2016; Changpinyo et al., 2021) to train OVS models under weak supervision (Xu et al., 2022a; Luo et al., 2023; Xu et al., 2023b; Zhang et al., 2023; Mukhoti et al., 2023; Cha et al., 2023; Liu et al., 2024a).
+The first paradigm offers many heuristic solutions based on observations of CLIP without additional training [81, 59, 33, 55, 29, 30].
+The second paradigm utilizes image-text pair data [57, 7] to train OVS models under weak supervision [69, 40, 70, 78, 44, 6, 39].
 Due to the lack of precise segmentation locations, these methods suffer from poor segmentation quality.
-The third paradigm enhances segmentation quality by incorporating image-mask data (Kirillov et al., 2023) and knowledge from SAM (Wang et al., 2024c; Yuan et al., 2024).
+The third paradigm enhances segmentation quality by incorporating image-mask data [28] and knowledge from SAM [63, 77].
 However, the use of SAM not only increases inference cost but also relies on users or other detectors to provide high-quality prompts.
 The fourth paradigm seeks to achieve mutual adaptation between the segmentation model and CLIP through fully supervised training on image segmentation datasets containing image, mask, and category triplets, thereby producing an open-vocabulary segmentation model.
-Based on the type of the segmentation model, these methods can be categorized into mask-based approaches (Ding et al., 2022; Xu et al., 2022b; Ghiasi et al., 2022; Han et al., 2023; Liang et al., 2023; Qin et al., 2023; Jiao et al., 2023; Ding et al., 2023; Xu et al., 2023d; c; a; Yu et al., 2024; Liu et al., 2024b; Jiao et al., 2024) and pixel/patch-based ones (Li et al., 2022; Zhou et al., 2023; Li et al., 2024; Cho et al., 2024; Xie et al., 2024).
+Based on the type of the segmentation model, these methods can be categorized into mask-based approaches [14, 72, 19, 20, 35, 46, 26, 15, 74, 73, 71, 76, 38, 27] and pixel/patch-based ones [31, 82, 34, 11, 68].
 Pixel-based OVS methods typically focus on semantic segmentation and cannot perform instance-level segmentation independently.
-Additionally, some unique approaches to achieving OVS involve image retrieval (Shin et al., 2022; Barsellotti et al., 2024), feature distillation (Chen et al., 2023; Wu et al., 2024) or model merging (Wang et al., 2024b).
-Some unified large models also incorporate the functionality of OVS (Zou et al., 2023; Shen et al., 2024).
+Additionally, some unique approaches to achieving OVS involve image retrieval [52, 1], feature distillation [8, 65] or model merging [61].
+Some unified large models also incorporate the functionality of OVS [83, 51].
 This paper focuses on efficiently transferring CLIP on limited segmentation datasets to improve its performance on semantic- and instance-level open-vocabulary segmentation tasks.
-3 Methodology
+3 Method
 3.1 Revisiting Image-Text Embeddings Alignment for OVS
@@
 In this space, image-text pairs maintain semantic alignment, which we describe as “pre-alignment” and can often be measured using cosine similarity $\langle E_{i},E_{t}\rangle$.
-In open-vocabulary segmentation, preserving the pre-alignment property is crucial for adapting CLIP to zero-shot dense predictions, as it enhances the alignment of region-specific image and text embeddings.
+In open-vocabulary segmentation, preserving the pre-alignment property is crucial for adapting CLIP to zero-shot mask classification, as it enhances the alignment of mask and text embeddings.
 Previous mask-based OVS methods obtain pre-aligned mask embeddings and text embeddings in two ways.
-As illustrated in Figure 2(a), the first way aims to train a mask generator that adapts to CLIP, with the generalization mainly derived from the original frozen parameters of CLIP (Xu et al., 2023c; Ding et al., 2023).
-However, attempts to align the embedding spaces of the mask generator and CLIP often fail (Xu et al., 2022b; Yu et al., 2024).
-As illustrated in Figure 2(b), the second way focuses on making CLIP sensitive to the generated masks during training to enhance performance and avoids overfitting through distillation (Jiao et al., 2023; 2024).
+As illustrated in Fig.
+2(a), the first way aims to train a mask generator that adapts to CLIP, with the generalization mainly derived from the original frozen parameters of CLIP [73, 15].
+However, attempts to align the embedding spaces of the mask generator and CLIP often fail [72, 76].
+As illustrated in Fig.
+2(b), the second way make CLIP sensitive to the quality of generated masks during training to enhance performance and avoids overfitting through distillation [26, 27].
 We identify two limitations of these two types of methods.
-First, they rely on noisy generated masks, which can degrade the quality of the region representations (as shown by the vertical comparisons in Figure 1(a)).
+First, they rely on noisy generated masks, which can degrade the quality of the region representations (as shown by the vertical comparisons in Fig.
+1(b)).
 This degradation can weaken the alignment between region-specific image and text embeddings.
@@
 To overcome these limitations, we propose a new training pipeline for OVS, named MaskCLIP++, which, to our knowledge, is the first method that attempts to leverage ground truth (GT) masks to obtain region features instead of relying on the mask generator during training.
-Furthermore, we propose the constraint named “consistency alignment” to alleviate overfitting when adding extra parameters to optimize alignment.
+Furthermore, we propose the principle named “consistency alignment” to alleviate overfitting when adding extra parameters to optimize alignment.
 We will give detailed descriptions in what follows.
-(a) Adapt Mask Generator to CLIP
-(b) Adapt CLIP to Mask Generator
-(c) MaskCLIP++
-Figure 2: Comparison of training pipeline between previous mask-based OVS methods.
-$\mathcal{L}$ denotes the loss function.
-(2(a)) adapts the mask generator to CLIP and avoids overfitting by freezing CLIP.
-(2(b)) adapts CLIP to the mask generator and avoids overfitting by distillation.
-“T” and “S” denote the teacher and student models, respectively.
-Our method (2(c)) abandons the mask generator and avoids overfitting by consistency alignment.
+Figure 3: Detailed framework of MaskCLIP++ for OVS tasks.
+The PSM represents the parameterized similarity modeling, which is designed under the principle of consistency alignment.
+The mask generator is only used during inference, and can be flexibly replaced.
 3.2 A New Fine-tuning Framework for CLIP
-In this subsection, we detail how to use our framework to fine-tune CLIP and integrate it with an off-the-shelf mask generator. The overall framework is illustrated in Figure 3.
+In this subsection, we detail how to use our framework to fine-tune CLIP and integrate it with an off-the-shelf mask generator. The overall framework is illustrated in Fig. 3.
 Extracting mask embeddings.
-Disregarding architectural details, CLIP-V can be considered as consisting of two parts: feature extraction by the encoder and feature aggregation by the neck.
-The mask is inserted after the last feature $F_{n}\in\mathbb{R}^{C\times H\times W}$ of the encoder is obtained, converting the original global aggregation into an aggregation within the mask’s effective region as follows:
-$E_{m}=\operatorname{Proj}\sum_{i,j}\left({\phi(M)}_{i,j}\cdot{(F_{n})}_{:,i,j}\right),$ (1)
-where $M$ is an object mask that has been downsampled to the same size as $F_{n}$, $E_{m}\in\mathbb{R}^{1\times D}$ is the embedding representations of this mask, $i,j$ are the spatial coordinates, and $\operatorname{Proj}$ represents the formulation of the neck that implements channel projection.
-The weighting function $\phi$ is related to the original CLIP pooling structure.
-Specifically, if the original CLIP uses global average pooling, $\phi$ acts as an L1 normalization function to achieve mask average pooling.
-If the original CLIP uses attention pooling, $\phi$ can be expressed as follows:
-$\phi(M)=\operatorname{Softmax}\left(\frac{q(T_{\text{cls}})k(T_{ij})}{\sqrt{d}}+\alpha M_{\text{bin}}\right),$ (2)
-where $T_{\text{cls}}$ and $T_{ij}$ represent the global and local representations sent to the attention pooling, respectively.
-$q,k$ are the input projection of the attention.
-$d$ is the hidden size of attention.
-The spatially expanded and binarized version of $M$ is denoted as $M_{\text{bin}}$, where any elements less than $\max(M)/2$ are set to $\infty$- ∞.
+Taking CLIP-V with an $L$-layer Transformers as an example, inspired by SAN [73], we posit that the layers from $1$ to $K$ predominantly perform feature extraction (denoted as the Extractor), while layers from $K+1$ to $L$ primarily conduct spatial fusion (referred to as the Fuser).
+For the $N$ image tokens $F^{(l)}\in\mathbb{R}^{N\times C}$ that input to the $l$-th layer of the fusion module, we introduce $Q$ flattened image masks $M\in[0,1]^{Q\times N}$ as conditions to transform the fusion scope from global to local.
+As formally delineated in Eq.
+1, we compute mask-conditioned modulation weights $\phi^{l}$ to perform weighted aggregation on $F^{(l)}$, followed by the layer-wise channel projection operator $\mathrm{Proj}^{(l)}$, thereby generating mask-specific tokens $E_{m}^{(l+1)}\in\mathbb{R}^{Q\times C}$.
+$E_{m}^{(l+1)}=\operatorname{Proj}^{(l)}\left(\phi^{(l)}(M)\cdot F^{(l)}\right),~{}\text{where}~{}K<l<L.$ (1)
+By leveraging the input projections $q$ and $k$ from the original attention module, we incorporate the mask into the weighting function through Eq.
+2, thereby constraining the weight distribution within the mask’s active regions.
+$\phi^{(l)}(M)=\operatorname{softmax}\left(\frac{q(E_{m}^{(l)})k(F^{(l)})^{T}}{\sqrt{d}}+\alpha M_{\text{th}}\right),$ (2)
+where $\sqrt{d}$ is the attention scaling factor, $\alpha\in(0,+\infty)$ is an learnable parameter, $M_{\text{th}}$ is the result of setting all values in $M$ that are less than $\max(M)/2$ to $\infty$- ∞.
+When initializing $E_{m}^{(K+1)}$ with the CLS token in layer $K+1$, $E_{m}^{(L)}$ and $E_{t}$ exhibit pre-aligned properties.
+The first two rows of Tab.
+1 quantitatively demonstrate the mask classification performance of pre-trained CLIP under this initialization paradigm.
+Fig.
+4 visualizes the normalized weight function $\phi$, which reveals the contributions of image tokens at different spatial positions to the generation of the current mask representation.
+After fine-tuning, more image tokens are incorporated into the construction of their corresponding mask representations.
+Figure 4: Visualization of the weight function. The regions of interest are enclosed by yellow contours, and the deeper the red color at a position, the higher its importance to the region.
 Parameterized similarity modeling with consistency alignment.
-The pretrained CLIP mask embedding $E_{m}$ and the category text embedding $E_{t}$ are initially pre-aligned with the initial similarity map $S=\langle E_{m},E_{t}\rangle$ being relatively coarse.
-Optimizing this alignment with some supervision is needed to achieve a more accurate similarity map.
-We refer to this process as parameterized similarity modeling (PSM).
-However, when incorporating the mask generator into similarity modeling, the mask embeddings produced by the mask generator align almost exclusively with CLIP’s text embeddings on the training set, resulting in the misalignment illustrated in Figure 2(a).
+Given the limited capacity of the original CLIP model in characterizing localized regions, a straightforward approach involves fine-tuning the original CLIP parameters alongside additional learnable parameters to reformulate the similarity metric between $E_{m}$ and $E_{t}$.
+We introduce a class of methods termed Parameterized similarity modeling (PSM) for this purpose.
+However, when incorporating the mask generator into similarity modeling, the mask embeddings produced by the mask generator align almost exclusively with CLIP’s text embeddings on the training set, resulting in the misalignment illustrated in Fig.
+2(a).
 We first analyze the cause of this overfitting to develop our similarity modeling approach.
@@
 Although $r$ can be optimized to be close to the target, an unintuitive phenomenon may arise: After mask embeddings are updated via gradient descent, the order of $r_{1}$ and $r_{2}$ is closer to the target, but the order of $s_{1}$ and $s_{2}$ may deviate further from the target.
-This phenomenon is illustrated by a toy example in Appendix Figure 5.
+This phenomenon is illustrated by a toy example in Sec.
+A.2 in supplementary material.
 Consequently, the original similarity relationships between CLIP’s embeddings are disrupted.
 Our experiments demonstrate that the additional parameter $\Theta$ fails to achieve CLIP’s level of generalization when fine-tuning on limited image segmentation data, resulting in overfitting.
-Therefore, when attempting to parameterize similarity modeling, it is essential to ensure that the modeled similarity maintains the same order as the original similarity, so that both can be optimized toward the same target.
-We refer to this constraint as “consistency alignment”.
-To optimize under the consistency alignment constraint in most of the time, the simplest way to achieve this is to update the similarity map in a dimension that is irrelevant to any modality using linear layers.
-Specifically, for $Q$ masks and $K$ texts, the $Q\times K$ similarity matrix $S$ is first expanded to $Q\times K\times T$ using a linear layer of size $T$, and then it is reduced to $Q\times K$ using another linear layer of the same size.
-When using the image patch embeddings $E_{p}$ to construct pseudo text embeddings (Jiao et al., 2024), a residual connection to the original text embeddings is necessary to ensure consistency alignment to some extent, as shown below:
-$S=\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle,\quad\text{where}~{}P_{t}=\operatorname{Attention}(q(E_{t}),k(E_{p}),v(E_{p}))$ (3)
-Here, $\operatorname{Attention}$ denotes the scaled dot-product attention (Vaswani, 2017), $q,k,v$ denote the input projection functions for attention and $\operatorname{Norm}$ denotes the layer norm (Lei Ba et al., 2016).
-The image patch embeddings $E_{p}$ are equivalent to specialized mask embeddings extracted from unit masks on a per-patch basis.
-Figure 3: The detailed framework of MaskCLIP++.
-The PSM represents the parameterized similarity modeling, which is constrained by the consistency alignment.
-The mask generator is only used during inference, and can be flexibly replaced.
+Building upon the aforementioned analysis, we identify inconsistent alignment phenomena in PSM that prone to overfitting.
+This implies that preserving CLIP’s original similarity ordering to the greatest extent possible during optimization enables more effective utilization of additional parameters for alignment enhancement.
+We formalize this critical finding as the consistency alignment principle in PSM.
+Specifically, for all mask embeddings $E_{m}\in\mathbb{R}^{Q\times C}$ extracted from an image and the text embeddings $E_{t}\in\mathbb{R}^{K\times C}$ of $K$ dataset categories, their similarity matrix $S=E_{m}E_{t}^{T}$ was projected through an over-parameterized linear projection along a newly created orthogonal dimension relative to both $Q$ and $K$ dimensions.
+Finally, cross-entropy loss was computed based on category annotations of each mask, guided by the optimized similarity map.
 Integrating with off-the-shelf mask generators.
 As the training process is entirely decoupled from the mask generator, we can flexibly employ various types of mask generators during inference to achieve semantic- or instance-level open-vocabulary segmentation.
-Moreover, integration with the class probabilities of the mask generator can appropriately improve the mask classification performance in the training categories.
+Compared to previous OVS methods, MaskCLIP++ offers the following several advantages.
+First, we use ground-truth masks to directly fine-tune CLIP during training, which are less noisy, thus improving alignment quality and reducing training cost.
+Second, from the perspective of avoiding overfitting, the consistency alignment principle provides more optimization space than freezing CLIP [73, 76], and does not rely on carefully designed distillation strategies [65, 27], resulting in faster convergence with a single optimization objective.
+4 Experiments
+4.1 Experimental Setup
+Baseline.
+In this section, we employ pre-trained models from EVA02 CLIP [54] with ViT architecture.
+Following the framework outlined in Sec.
+3.2, we configure $L-K=2$ with $\alpha$ initialized to $e^{-5}$.
+The second row of Tab.
+1 demonstrates the mask classification accuracy of this baseline implementation.
+For open-vocabulary image segmentation evaluation, we adopt FC-CLIP (corresponding to Fig.
+2(a)) and MAFT+ (corresponding to Fig.
+2(b)) as baseline approaches.
+Their baseline performance metrics are documented in Tab.
+2 and Tab.
+3.
+By retaining the mask generators from both frameworks while replacing the mask classification component with MaskCLIP++, we demonstrate the performance enhancement in overall OVS capabilities through improved mask classification accuracy.
+Datasets and Metrics.
+Aligning with established OVS methodologies, we fine-tune our model on the COCO Stuff dataset [5] with its fixed 171-category vocabulary and evaluate on datasets with different vocabularies: ADE20K (150-class and 847-class taxonomies) [80], Pascal Context (459-class and 59-class taxonomies) [43], and PASCAL VOC (20 foreground categories excluding background) [17].
+The mean Intersection over Union (mIoU) metric serves as our primary evaluation criterion for OVS performance.
+Furthermore, we quantitatively demonstrate performance enhancements in both panoptic and instance segmentation tasks on ADE20K through established metrics: panoptic quality (PQ), segmentation quality (SQ), recognition quality (RQ) and mean average precision (mAP).
+To demonstrate the generalizable improvement of MaskCLIP++ in mask classification in various scenarios, we perform evaluations on datasets from the MESS benchmark [4], where the accuracy of the top-1 ground truth mask (maskAcc) serves as the primary metric to quantify the capability of zero shot mask classification.
+Implementation Details.
+Our implementation is built on the Detectron2 framework [67].
+During training, the number of image tokens $N=32^{2}$.
+The AdamW optimizer is employed with a batch size of 4 for 20K iterations.
+The initial learning rate is set to $10^{-4}$ and decayed using cosine annealing.
+Following the setting of CAT-Seg [11], we only train the linear projection parameters of the query and value in the CLIP attention modules, and set their learning rate $10^{2}$ times the overall learning rate.
+The over-parameterized dimension of the linear layer in PSM is set to 768.
+During inference, while maintaining the original aspect ratio, images are resized with shorter edges scaled to 392 and 800 pixels respectively for CLIP-V and Mask Generator inputs.
+Similar to previous works [71, 76], we also ensemble the class scores of the mask generator and MaskCLIP++ on the training categories of the mask generator.
+A-847 PC-459 A-150 PC-59 Stuff Citys General Earth Medical Engineer Agriculture
+Pretrained
+OpenAI CLIP [47] 32.0 44.8 51.0 57.9 46.2 46.5 54.2 61.1 61.0 36.7 51.0
+EVA02 CLIP [54] 35.2 44.8 52.7 54.6 45.0 44.9 56.9 60.5 61.7 33.8 52.4
+Finetuned on COCO
+CLIPSelf [65] 33.6 49.0 56.1 67.5 50.6 52.1 52.5 45.7 55.5 45.3 56.3
+CAT-Seg [11] 33.5 50.7 62.3 80.6 63.0 65.3 60.7 66.9 62.0 34.9 64.4
+MaskCLIP++ 38.4 56.4 67.0 85.2 67.8 71.0 67.9 68.6 74.7 50.3 65.5
+Table 1: Comparison of the classification accuracy of ground-truth masks (maskAcc) with different methods on multiple datasets.
+The same architecture (ViT-L/14) and input resolution (short side length 392) are used for all methods.
+“General”, “Earth”, “Medical”, “Engineer”, and “Agriculture” represent mean mask accuracy across five domain datasets in the MESS benchmark [4], with detailed scores for each individual dataset presented in the supplementary material.
+Method CLIP arch. Training dataset A-847 PC-459 A-150 PC-59 PAS-20
+Non-mask-based OVS method
+OVSeg [35] ViT-L/14 COCO Stuff 9.0 12.4 29.6 55.7 94.5
+CLIPSelf [65] ViT-L/14 COCO Stuff 12.4 - 34.5 62.3 -
+SED [68] ConvNeXt-L COCO Stuff 13.7 22.1 35.2 60.6 96.1
+CAT-Seg [11] ViT-L/14 COCO Stuff 16.0 23.8 37.9 63.3 97.0
+Mask-based OVS method
+MaskCLIP [15] ViT-L/14 COCO Panoptic 8.2 10.0 23.7 45.9 -
+MasQCLIP [74] ViT-L/14 COCO Panoptic 10.7 18.2 30.4 57.8 -
+ODISE [71] ViT-L/14 COCO Panoptic 11.1 14.5 29.9 57.3 -
+SAN [73] ViT-L/14 COCO Stuff 12.4 15.7 32.1 57.7 94.6
+SCAN [38] ViT-L/14 COCO Stuff 14.0 16.7 33.5 59.3 97.2
+FC-CLIP [76] ConvNeXt-L COCO Panoptic 14.8 18.2 34.1 58.4 95.4
+MAFT+ [27] ConvNeXt-L COCO Stuff 15.1 21.6 36.1 59.4 96.5
+MaskCLIP++ w/ FC-CLIP ViT-L/14 COCO Stuff 15.4+0.6 21.3+3.1 37.1+3.0 62.6+4.2 96.4+1.0
+MaskCLIP++ w/ MAFT+ ViT-L/14 COCO Stuff 16.8+1.7 23.9+2.3 38.2+2.1 62.5+3.1 96.8+0.3
+Table 2: Comparisons with previous methods on open-vocabulary semantic segmentation task.
+4.2 Main results
+Mask classification.
+Tab.
+1 presents the mask classification accuracy of CLIP models from different sources when processing input images and ground-truth masks at identical resolutions.
+In addition to the two pre-trained CLIP variants, we also demonstrate the performance of CLIP fine-tuned on COCO through distinct approaches: CLIPSelf [65] learned via patch-level feature distillation and CAT-Seg [11] trained with pixel-level cost aggregation.
+During inference, CLIPSelf performs mask pooling on the dense features of the final layer.
+Given that CAT-Seg operates as a pixel-level segmentation model, we leverage its CLIP parameters while maintaining consistency with the baseline methods by employing identical mask classification methodology.
+By directly utilizing higher-quality masks to fine-tune CLIP while adhering to consistent alignment principles, MaskCLIP++ achieves significant improvements over baselines across multi-domain datasets, outperforming other fine-tuning approaches.
+Open-vocabulary semantic segmentation.
+Tab.
+2 demonstrates the mIoU of MaskCLIP++ when employing mask generators from FC-CLIP [76] or MAFT+ [26].
+Compared to MAFT+, our approach achieves mIoU improvements of +1.7, +2.3, +2.1, +3.2, and +0.3 points on the A-847, PC-459, A-150, PC-59, and PAS-20 datasets, respectively.
+These results indicate that our method more effectively harnesses the potential of existing mask generators through better mask classification capabilities.
+Open-vocabulary panoptic and instance segmentation.
+Tab.
+3 demonstrates that our method can be effectively integrated with existing instance-level mask generators.
+Furthermore, the enhanced classification capabilities of MaskCLIP++ also improve the performance across these segmentation paradigms.
+Method PQ SQ RQ AP
+MaskCLIP [15] 15.1 70.5 19.2 6.2
+MasQCLIP [74] 23.3 - - -
+ODISE [71] 22.6 - - 14.4
+FC-CLIP [76] 26.8 71.5 32.2 16.8
+MAFT+ [27] 27.1 73.5 32.9 -
+MaskCLIP++ w/ FC-CLIP 27.7 72.0 33.6 17.3
+MaskCLIP++ w/ MAFT+ 28.1 74.0 34.7 -
+Table 3: Comparisons with previous methods on open-vocabulary panoptic and instance segmentation on A-150.
+4.3 Ablation Studies
+The ablation studies in Tab.
+4 and Tab.
+5 are conducted using ViT-B/16 models fine-tuned on COCO Panoptic [36], whereas those presented in Tab.
+6 and Tab.
+7 employed ViT-L/14 models fine-tuned on COCO Stuff [5].
+Importance of consistency alignment.
+We use different parameterized similarity modeling methods to illustrate the importance of the consistent alignment principle.
+Tab.
+4 elucidates the relationship between consistency alignment and overfitting through PSM application analysis during inference.
+We evaluate alignment consistency by toggling PSM usage after fine-tuning, while partitioning validation set categories into seen (S) and unseen (U) subsets to compute mIoU(S) and mIoU(U) respectively.
+Severe overfitting is identified when optimized mIoU(U) significantly underperforms its no-fine-tuning baseline.
+The table substantiates that the $\operatorname{Linear}\langle E_{m},E_{t}\rangle$ approach preserving consistency alignment achieves simultaneous improvements in both mIoU(S) and mIoU(U), demonstrating effective generalization maintenance.
+PSM Use mIoU(S) mIoU(U) mIoU
+No fine-tune - 22.6 12.8 17.0
+$\langle\operatorname{Linear}(E_{m}),E_{t}\rangle$ 36.6 21.7 28.0
+40.4 6.2 20.8
+$\langle E_{m},\operatorname{Linear}(E_{t})\rangle$ 36.1 21.4 27.7
+41.2 6.6 21.4
+$\operatorname{Linear}\langle E_{m},E_{t}\rangle$ 44.8 25.1 33.5
+45.3 25.3 33.8
+Table 4: The phenomenon of (in)consistent alignment when fine-tuning with different PSMs.
+The PSMs prone to overfitting (red) paradoxically shows better performance when omitting the PSM during inference, revealing the occurrence of inconsistent alignment.
+In contrast, the blue PSM maintains consistent alignment while achieving comparable performance both with and without the PSM application, thus effectively mitigating overfitting risks.
+Impact of prior quality during training.
+We refer to the regional information fed into Fuser of CLIP-V as “priors” and investigate how the quality of these priors affects fine-tuning performance.
+We experiment with pixels, generated masks that match GT masks according to IoU, GT masks, and the GT boxes as their minimum bounding rectangles.
+As shown in Tab.
+5, using GT masks can achieve the best performance.
+Using the generated masks as prior yields lower performance compared to the GT masks, indicating that incorporating the mask generator during training actually hinders CLIP fine-tuning.
+Both GT masks and GT boxes priors outperform pixel priors, suggesting that fine-tuning CLIP from global recognition to region-level recognition is less challenging than fine-tuning it to pixel-level recognition.
+Prior Type mIoU(S) mIoU(U) mIoU maskAcc
+Pixels 43.2 21.4 30.6 53.3
+Gen. masks 43.8 23.8 32.3 52.2
+GT boxes 44.4 24.2 32.8 57.6
+GT masks 45.3 25.3 33.8 58.3
+Table 5: Impact of prior quality during training. “Gen. masks” denotes the matched generated masks. “GT boxes” denotes the minimum bounding rectangles of GT masks.
+Impact of different mask generators.
+Except for FC-CLIP and MAFT+ in Tab.
+2, various types of mask generators can be integrated with MaskCLIP++, as shown in Tab.
+6.
+We use mask generators from Mask2Former [10] and FC-CLIP [76], which are typically considered as closed-vocabulary and open-vocabulary mask generators, respectively.
+These mask generators are all trained on COCO Panoptic [36] and evaluated on ADE20K [80] using mIoU.
+As shown in Tab.
+6, the original performance of “open-vocabulary” mask generators do do not significantly outperform “closed-vocabulary” ones because of the “mis-alignment” as shown in Fig.
+2(a).
+And using our MaskCLIP++ for mask classification can significantly improve the segmentation performance of all mask generators.
+This suggests that both open-vocabulary and closed-vocabulary mask generators possess an inherent category-independent mask generation capability that has yet to be fully harnessed.
+Generator OV Gen. Only w/ MaskCLIP++
+Mask2Former (T) 15.7 35.6
+Mask2Former (B) 16.8 35.5
+FC-CLIP (B) 19.1 36.8
+FC-CLIP (L) 20.8 37.1
+Table 6: Impact of different mask generators on OVS performance.
+“OV” indicates whether the mask generator is open-vocabulary.
+“Gen.
+Only” indicates the mIoU of the mask generator on A-150.
+“w/ MaskCLIP++” indicates the mIoU of the mask generator when integrated with MaskCLIP++.
+Impact of training data size.
+Since the goal of our fine-tuning is to transfer CLIP’s open-vocabulary recognition capabilities to mask regions, rather than learning category-specific pixel groupings from segmentation datasets, our method can effectively fine-tune using a smaller amount of image segmentation data.
+We randomly sampled varying proportions of data from the COCO-Stuff training set for fine-tuning, using the same configuration as the model in Tab.
+2.
+As shown in Tab.
+7, even with only 1% of the data (approximately 1K images with annotations), the fine-tuned model still demonstrates competitive performance.
+Data size A-847 PC-459 A-150 PC-59 Stuff
+100% 38.4 56.4 67.0 85.2 67.8
+10% 39.4 56.2 67.0 84.0 67.4
+1% 38.7 55.4 66.3 83.5 65.5
+0.1% 36.3 54.5 61.2 75.5 55.7
+Table 7: Mask accuracy when fine-tuning with different amounts of data. “data size” indicates the proportion of data used for fine-tuning relative to the original COCO-Stuff training set.
+Image                           CAT-Seg                        MAFT+                     MaskCLIP++ GT
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Figure 5: Visualizations of open-vocabulary semantic segmentation on ADE20K [80]. Our method produces more complete masks than CAT-Seg and fewer biases in mask classification than MAFT+.
+5 Discussion
+Qualitative comparison.
+Fig.
+5 presents qualitative comparisons of CAT-Seg [11], MAFT+ [27], and MaskCLIP++ on open-vocabulary semantic segmentation tasks.
+Compared to CAT-Seg, mask-based methods often produce higher-quality masks.
+Compared to MAFT+, our method retains more of the original CLIP’s knowledge without generating too many erroneous biases.
+Our method also predicts relatively correct semantics for some objects that are not annotated in the GT.
+Segmentation without mask generators.
+Because the goal of our fine-tuning is to enhance the mask classification ability of CLIP, which indirectly enhances the local representation between different patch representations of the CLIP.
+Diverging from Sec.
+3.2, we apply mask generator-free inference schemes on MaskCLIP++ to demonstrate the enhanced inter-patch relationships, as shown in Tab.
+8.
+Method Load PC-459 A-150 PC-59 Stuff Citys
+SCLIP [59] 7.7 16.1 34.2 22.4 32.2
+9.5 19.3 40.8 29.0 35.9
+ClearCLIP [29] 7.9 16.7 36.0 23.9 30.0
+9.6 19.2 40.8 30.4 33.5
+ProxyCLIP  [30] 8.4 20.2 39.1 26.5 38.1
+9.9 23.2 44.6 34.2 42.4
+Table 8: Performance improvements achieved when combined with several unsupervised open-vocabulary semantic segmentation methods that do not rely on a mask generator.
+“Load” means whether to load model weights from MaskCLIP++.
+OpenAI CLIP ViT-B/16 are used here.
+Additional computation beyond mask generators.
+Tab.
+9 presents the computational costs of different modules in our model from Tab.
+2 under varying numbers of proposal masks.
+The CLIP-V adopts the ViT-L/14 architecture with an input resolution of $392^{2}$, while the mask generator employs the ConvNeXt-L architecture operating at $800^{2}$ resolution.
+Using 150 object categories, the table reveals that the primary computational bottleneck remains in the mask generator since CLIP-V for mask classification typically requires lower resolutions.
+Furthermore, the Extractor and mask generator can execute computations in parallel during implementation, thereby reducing overall processing time.
+Q CLIP-V PSM Mask Generator
+Extractor Fuser Backbone Decoder
+100 416.26 67.05 0.05 877.63 232.12
+250 416.26 78.62 0.12 877.63 302.35
+Table 9: FLOPs (G) of each module.
+6 Conclusion
+This paper addresses key challenges in open-vocabulary segmentation (OVS) by proposing MaskCLIP++, a high-quality mask tuning framework for CLIP that significantly simplifies the training paradigm.
+Our method achieves notable improvements in mask classification accuracy across multiple domain datasets, and when integrated with existing mask generators, demonstrates superior open-vocabulary image segmentation performances.
+Supplementary Material for “High-Quality Mask Tuning Matters for Open-Vocabulary Segmentation”
+Appendix A Method Details
+A.1 Extracting mask embeddings of different architectures
+The main text exclusively details mask embedding extraction from ViT-architecture CLIP models.
+This methodology can be extended to other CLIP architectures through the following adaptations: For architectures containing attention pooling layers without Transformer blocks, the Fuser module operates starting from the attention pooling layer, with weight functions identical to Eq.
+2.
+In models employing global average pooling without Transformer components, the Fuser initiates computation from the global pooling layer, where the weight function degenerates to L1 normalization to implement masked average pooling.
+A.2 Phenomenon of inconsistency alignment
+We use a toy example to illustrate the potential consequences of destroying alignment consistency.
+Assume that we use the parameter $\Theta$ to update the mask embeddings, changing the original similarity from $s$ to $r$, and then optimize towards the target.
+However, due to the lack of explicit constraints on $\Theta$, there may be a situation as shown in Fig.
+6, where the order of $s$ and $r$ is inconsistent, resulting in correct optimization on $r$ but incorrect optimization on $s$.
+For simplicity, $\mathcal{L}$ in Fig.
+6 denotes an L1 loss without softmax, and only mask embeddings $m$ are updated in one round of gradient descent.
+A.3 Mask classification score integration
+The mask generator demonstrates inherent capability to produce classifications within its training vocabulary.
+When unambiguous semantic correspondence exists during inference, following established protocols in mask-based open-vocabulary segmentation (OVS) works [72, 71, 76], our method’s output class scores can be adaptively aggregated across the mask generator’s training vocabulary.
+This enables category probability estimation while maintaining compatibility with existing OVS frameworks.
 Specifically, let $\mathcal{K}$ be the training vocabularies of the mask generator.
-$P_{c}$ and $P_{s}$ are the class probabilities of the fine-tuned CLIP and the mask generator, respectively.
+$P_{c}$ and $P_{s}$ are the class probabilities of /Ours and the mask generator, respectively.
 The integrated class probabilities $P_{\gamma}$ can be expressed as follows:
 $P_{\gamma}^{(i)}=\begin{cases}{P_{s}^{(i)}}^{\gamma}\cdot{P_{c}^{(i)}}^{(1-\gamma)},&\text{if}~{}i\in\mathcal{K}\\ P_{c}^{(i)}.&\text{otherwise}\\ \end{cases}$ (3)
-Compared to previous OVS methods, our MaskCLIP++ offers the following several advantages.
-First, we use GT masks during training, which are less noisy and more accurate, thus improving alignment quality and reducing training cost.
-Second, from the perspective of avoiding overfitting, the consistency alignment constraint provides more optimization space than freezing CLIP (Xu et al., 2023c; Yu et al., 2024), and does not rely on carefully designed distillation strategies (Wu et al., 2024; Jiao et al., 2024), resulting in faster convergence with a single optimization objective.
-4 Experiments
-4.1 Implementation Details
-In the following experiments, we utilize various architectures of pre-trained CLIP to demonstrate the generality of our method.
-The CLIP models with ResNet architecture are sourced from OpenAI (Radford et al., 2021), those with ViT architecture from EVA-CLIP (Sun et al., 2023), and those with ConvNeXt architecture from OpenCLIP (Cherti et al., 2023).
-By default, consistent with previous work on universal open-vocabulary semantic segmentation (Yu et al., 2024), we fine-tune our model using the COCO Panoptic training set (Lin et al., 2014), which consists of 118K images and 133 categories.
-The AdamW optimizer is utilized with an initial learning rate of 2e-4, decaying to 1e-5 via cosine annealing.
-The fine-tuning process is conducted for 20K iterations with a batch size of 4.
-The size of the training image is adjusted according to the downsampling factor of the model to ensure that the final size of the feature map is $32^{2}$.
-The parameters involved in fine-tuning differ slightly across architectures.
-For the ConvNeXt and ResNet architectures, all parameters of CLIP-V are fine-tuned.
-For the ViT architecture, following Cho et al.
-(2024), only the positional encodings and embedding layers of CLIP-V are fine-tuned.
-In the ConvNeXt architecture, the fine-tuning learning rate for CLIP parameters is 1e-3 times that of other parts, while in the ResNet and ViT architectures, it is 1e-2 times.
-During inference, we mainly evaluate using the ADE20K validation set (Zhou et al., 2017), which contains 2,000 images and 150 categories, along with annotations for semantic segmentation, panoptic segmentation, and instance segmentation.
-In ADE20K, 64 categories overlap with those in COCO, while the remaining 86 categories do not.
-We measure semantic segmentation performance using Intersection over Union (IoU), recording mIoU(S) for overlapping categories, mIoU(U) for non-overlapping categories, and overall mIoU.
-Panoptic Quality (PQ) and Average Precision (AP) are used to assess performance in panoptic segmentation and instance segmentation, respectively.
-By default, we use FC-CLIP (Yu et al., 2024) trained on the COCO Panoptic (Lin et al., 2014) training set as the mask generator.
-For the base and large level CLIP, we employ the corresponding FC-CLIP models as mask generators  [^1] Since FC-CLIP did not provide a mask generator with CLIP ConvNeXt-B as the backbone, we used their code for CLIP ConvNeXt-L to produce a base one..
-The inference size of the mask generator remains consistent with FC-CLIP, set to 800.
-The inference size of CLIP depends on the model architecture, and the input sizes and initial segmentation performance for different architectures can be seen in Table 4.
-[…]
-PSM C.A. mIoU(S) mIoU(U) mIoU
-Baseline - 32.7 16.3 23.3
-$\langle\operatorname{Linear}(E_{m}),E_{t}\rangle$ 32.4 1.9 14.9
-$\langle E_{m},\operatorname{Linear}(E_{t})\rangle$ 32.5 1.8 14.9
-$\operatorname{Linear}\langle E_{m},E_{t}\rangle$ 45.4 25.7 34.1
-$\langle E_{m},\operatorname{Norm}(P_{t})\rangle$ 31.9 4.5 16.2
-$\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle$ 46.3 26.7 35.1
-Table 2: Impact of prior quality during training. “Seg masks” is the matched generated masks. “GT boxes” is the minimum bounding rectangles of GT masks.
-Prior Type mIoU(S) mIoU(U) mIoU
-Pixels 44.3 24.2 32.8
-Seg masks 44.9 23.7 32.7
-GT boxes 45.4 25.8 34.2
-GT masks 46.3 26.7 35.1
-4.2 Ablation Study of MaskCLIP++
-In our ablation study, we used CLIP ConvNeXt-B along with the mask generator from FC-CLIP (Yu et al., 2024), which shares the same backbone.
-The direct combination of them achieved a baseline mIoU of 23.3 on ADE20K (Zhou et al., 2017).
-Importance of consistency alignment.
-We use different parameterized similarity modeling methods to illustrate the importance of the consistent alignment constraint.
-As shown in Table 2, the attempts to use linear parameters to update mask embeddings or text embeddings violate the constraint of consistency alignment during fine-tuning, resulting in a significant performance drop on unseen categories compared to the baseline (1.9% vs.
-16.3% or 1.8% vs.
-16.3%).
-The $P_{t}$ are weighted sums of image patch embeddings as shown in equation 3, so $\langle E_{m},\operatorname{Norm}(P_{t})\rangle$ cannot have a pre-aligned initial similarity map, leading to substantial performance drops on unseen categories as well (4.5% vs.
-16.3%).
-The PSM of $\operatorname{Linear}\langle E_{m},E_{t}\rangle$ adheres to the consistency alignment constraint, resulting in performance improvements on both seen and unseen categories (45.4% vs.
-32.7% and 25.7% vs.
-16.3%).
-For $\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle$, the image context information $P_{t}$ indirectly interacts with $E_{m}$.
-This compensates for the rigidity of mask average pooling in the CLIP ConvNeXt architecture, ultimately resulting in improved performance (35.1% vs.
-34.1%).
-When fine-tuning CLIP models with attention pooling (ViT and ResNet), we still use the PSM of $\operatorname{Linear}\langle E_{m},E_{t}\rangle$.
-We present additional experiments in Appendix Table 12 to demonstrate the extent to evaluate how the PSMs in Table 2 preserve or disrupt the consistency alignment constraint.
-Impact of prior quality during training.
-We refer to the regional information fed into CLIP during the mask embedding extraction stage as “priors” and investigate how the quality of these priors affects fine-tuning performance.
-We experiment with pixels, generated masks that match GT masks according to IoU, GT masks, and the GT boxes as their minimum bounding rectangles.
-As shown in Table 2, using GT masks can achieve the best performance.
-Using the generated masks as prior yields lower performance compared to the GT masks (32.7% vs.
-35.1%), indicating that incorporating the mask generator during training actually hinders CLIP fine-tuning.
-Both GT masks and GT boxes priors outperform pixel priors, suggesting that fine-tuning CLIP from global recognition to region-level recognition is less challenging than fine-tuning it to pixel-level recognition.
-Table 3: Impact of different mask generators.
-“OV” indicates that the mask generator utilizes a pre-trained CLIP as the backbone and employs CLIP text embeddings as a variable-length classification head, typically considered an open-vocabulary mask generator.
-$P_{s}$, $P_{c}$ and $P_{\gamma}$ are defined in equation 4.
-mIoU are reported.
-Latency (ms) is the average inference time of a single image on ADE20K, measured in an environment with Pytorch 2.3.1, cuda 11.8 and one RTX 2080Ti GPU.
-Mask Generator + MaskCLIP++
-Backbone Decoder OV $P_{s}$ $P_{c}$ $P_{\gamma=0.1}$ $P_{\gamma=0.2}$ $P_{\gamma=0.4}$ Latency (ms)
-Swin-T Mask2Former 15.7 33.9 34.3 34.2 33.5 15.5
-Swin-L Mask2Former 16.8 33.9 34.3 33.9 33.2 33.6
-ConvNeXt-B FC-CLIP 19.1 35.1 35.7 35.8 35.0 32.3
-ConvNeXt-L FC-CLIP 20.8 34.6 35.4 35.8 35.8 39.0
-Table 4: Improvements of various CLIP architectures.
-All models are fine-tuned on COCO Panoptic and evaluated on ADE20K.
-“Res.” denotes the input resolution of CLIP during inference.
-For CNN-based CLIP, all parameters in CLIP-V are fine-tuned.
-For ViT-based CLIP, only the position embedding and input projections of the attention in the CLIP-V are fine-tuned.
-For RN50x16, ConvNeXt-B, and ViT-B/16, use the mask generator with ConvNeXt-B backbone.
-For others, use the mask generator with ConvNeXt-L backbone.
-Ensemble strategy is not used.
-CLIP Fine-tuned Input Original CLIP MaskCLIP++
-Arch. Part Params(M) Res mIoU PQ AP mIoU PQ AP
-RN50x16 V-All 167.3 512 18.4 15.8 9.4 29.3+10.9 21.8+6.0 11.1+1.7
-ConvNeXt-B V-All 90.3 512 23.3 19.1 11.6 35.1+11.8 24.5+5.4 13.6+2.0
-ConvNeXt-L V-All 202.7 640 25.4 20.7 14.6 35.6+10.2 26.5+5.8 16.7+2.1
-ConvNeXt-XXL V-All 846.5 640 26.3 20.3 13.3 36.4+10.1 27.1+6.8 16.6+3.3
-ViT-B/16 V-Attn 14.3 384 17.0 15.0 9.5 33.8+16.8 24.4+9.4 13.2+3.7
-ViT-L/14 V-Attn 50.9 392 20.3 17.4 12.4 36.6+16.3 27.3+9.9 17.0+4.6
-ViT-G/14 V-Attn 238.3 336 20.7 18.6 13.5 36.8+16.1 27.7+9.1 17.1+3.6
-Impact of different mask generators.
-Various types of mask generators can be integrated with our fine-tuned CLIP, as shown in Table 3.
-We use mask generators from Mask2Former (Cheng et al., 2022) and FC-CLIP (Yu et al., 2024), which are typically considered as closed-vocabulary and open-vocabulary mask generators, respectively.
-These mask generators are all trained on COCO Panoptic (Lin et al., 2014) and evaluated on ADE20K (Zhou et al., 2017) using mIoU.
-As shown in Table 3, open-vocabulary mask generators do not significantly outperform closed-vocabulary mask generators in terms of segmentation performance ($P_{s}$).
-And using our fine-tuned CLIP for mask classification can significantly improve the segmentation performance ($P_{c}$) of all mask generators.
-This suggests that both open-vocabulary and closed-vocabulary mask generators possess an inherent category-independent mask generation capability that has yet to be fully harnessed.
-Furthermore, integrating the classification scores of the mask generator on training categories, as described in equation 4, can slightly enhance performance.
-For the hyperparameter $\gamma$, empirically, the worse the mask classification quality $P_{s}$ of the mask generator is, the smaller $\gamma$ can be set.
-4.3 Enhancement of Mask Classification by MaskCLIP++
-We apply our method to CLIP with various architectures to demonstrate its universality.
-The experimental settings are described in Section 4.1.
-Table 4 presents the types of CLIP architectures, the parts that were fine-tuned, and the corresponding number of fine-tuned parameters.
-Compared to the original CLIP, our method improved mIoU by 10.1%-16.8%, PQ by 5.4%-9.9% and AP by 1.7%-4.6% on ADE20K(Zhou et al., 2017) when using COCO Panoptic (Lin et al., 2014) for fine-tuning.
-Additionally, we demonstrate significant performance improvements of the original CLIP across different input resolutions, as shown in Figure 6 in Appendix C.
-Table 5: Comparisons with previous methods on Open-Vocabulary Semantic Segmentation task.
-Method VLM Training dataset A-847 PC-459 A-150 PC-59 PAS-20
-Non-mask-based OVS method
-SED (Xie et al., 2024) CLIP ConvNeXt-B COCO-Stuff 11.4 18.6 31.8 57.7 94.4
-CAT-Seg (Cho et al., 2024) CLIP ViT-B/16 COCO-Stuff 12.0 19.0 31.8 57.5 94.6
-Mask-based OVS method
-ZSSeg (Xu et al., 2022b) CLIP ViT-B/16 COCO-Stuff 7.0 - 20.5 47.7 -
-OpenSeg (Ghiasi et al., 2022) ALIGN EN-B7 COCO-Captions + Loc. Narr. 8.8 12.2 28.6 48.2 72.2
-DeOP (Han et al., 2023) CLIP ViT-B/16 COCO-Stuff-156 7.1 9.4 22.9 48.8 91.7
-FreeSeg (Qin et al., 2023) CLIP ViT-B/16 COCO-Panoptic - - 24.6 - -
-MAFT (Jiao et al., 2023) CLIP ViT-B/16 COCO-Stuff 10.1 12.8 29.1 53.5 90.0
-SAN (Xu et al., 2023c) CLIP ViT-B/16 COCO-Stuff 10.1 12.6 27.5 53.8 94.0
-SCAN (Liu et al., 2024b) CLIP ViT-B/16 COCO-Stuff 10.8 13.2 30.8 58.4 97.0
-FC-CLIP (Yu et al., 2024) CLIP ConvNeXt-B COCO-Panoptic 13.1 16.5 31.2 55.2 94.2
-MAFT+ (Jiao et al., 2024) CLIP ConvNeXt-B COCO-Stuff 13.8 16.2 34.6 57.5 95.4
-MaskCLIP++ CLIP ConvNeXt-B COCO-Stuff 14.5 18.7 35.4 59.1 95.8
-Non-mask-based OVS method
-OVSeg (Liang et al., 2023) CLIP ViT-L/14 COCO-Stuff 9.0 12.4 29.6 55.7 94.5
-CLIPSelf (Wu et al., 2024) CLIP ViT-L/14 COCO-Stuff 12.4 - 34.5 62.3 -
-SED (Xie et al., 2024) CLIP ConvNeXt-L COCO-Stuff 13.7 22.1 35.2 60.6 96.1
-CAT-Seg (Cho et al., 2024) CLIP ViT-L/14 COCO-Stuff 16.0 23.8 37.9 63.3 97.0
-Mask-based OVS method
-MaskCLIP (Ding et al., 2023) CLIP ViT-L/14 COCO-Panoptic 8.2 10.0 23.7 45.9 -
-MasQCLIP (Xu et al., 2023d) CLIP ViT-L/14 COCO-Panoptic 10.7 18.2 30.4 57.8 -
-ODISE (Xu et al., 2023a) CLIP ViT-L/14 COCO-Panoptic 11.1 14.5 29.9 57.3 -
-SAN (Xu et al., 2023c) CLIP ViT-L/14 COCO-Stuff 12.4 15.7 32.1 57.7 94.6
-SCAN (Liu et al., 2024b) CLIP ViT-L/14 COCO-Stuff 14.0 16.7 33.5 59.3 97.2
-FC-CLIP (Yu et al., 2024) CLIP ConvNeXt-L COCO-Panoptic 14.8 18.2 34.1 58.4 95.4
-MAFT+ (Jiao et al., 2024) CLIP ConvNeXt-L COCO-Stuff 15.1 21.6 36.1 59.4 96.5
-MaskCLIP++ CLIP ViT-L/14 COCO-Stuff 16.8 23.9 38.2 62.5 96.8
-4.4 Comparison with previous methods
-Open-vocabulary semantic segmentation.
-There are notable differences in the training settings compared to Section 4.1 to be consistent with most previous open vocabulary segmentation works.
-We fine-tune CLIP on COCO-Stuff (Caesar et al., 2018), which shares the same images as COCO Panoptic (Lin et al., 2014) but defines 171 categories.
-We report mIoU performance on the validation sets of the following datasets: ADE-20K (Zhou et al., 2017) with 847 classes (A-847) and 150 classes (A-150), PASCAL-Context (Mottaghi et al., 2014) with 459 classes (PC-459) and 59 classes (PC-59), and PASCAL-VOC (Everingham et al., 2010) with 20 foreground classes (PAS-20).
-For comparisons with previous works using base-level VLM models, we fine-tune CLIP ConvNeXt-B with previous settings.
-For comparisons with works using larger-level VLM models, we fine-tune CLIP ViT-L/14.
-Apart from the previous settings, we also fine-tune the input projections of the attention in the CLIP text encoder following CAT-Seg (Cho et al., 2024).
-We use the same mask generator as MAFT+ (Jiao et al., 2024) during inference and set $\gamma=0.4$ for CLIP ConvNeXt-B and $\gamma=0.1$ for CLIP ViT-L/14.
-Table 5 shows that our base model achieves the best performance on A847, A150 and PC-59, and our large model achieves the best performance on A-847, PC-459, and A-150.
-Table 6: Comparisons with previous methods on Open-Vocabulary Panoptic and Instance Segmentation tasks.
-All methods are trained on COCO Panoptic (Lin et al., 2014) and evaluated on ADE20K (Zhou et al., 2017).
-Method VLM PQ AP
-FreeSeg (Qin et al., 2023) CLIP ViT-B/16 16.3 6.5
-FC-CLIP (Yu et al., 2024) CLIP ConvNeXt-B 23.1 13.1
-MaskCLIP++ CLIP ConvNeXt-B 24.5 13.6
-MaskCLIP (Ding et al., 2023) CLIP ViT-L/14 15.1 6.2
-MasQCLIP (Xu et al., 2023d) CLIP ViT-L/14 23.3 -
-ODISE (Xu et al., 2023a) CLIP ViT-L/14 22.6 14.4
-FC-CLIP (Yu et al., 2024) CLIP ConvNeXt-L 26.8 16.8
-MAFT+ (Jiao et al., 2024) CLIP ConvNeXt-L 27.1 -
-MaskCLIP++ CLIP ViT-L/14 27.3 17.0
-Table 7: Comparison of CLIP’s mask classification ability in different methods.
-The same ViT-L/14 architecture of CLIP and the same mask generator are used.
-The first three rows are initialized by OpenAI-CLIP, and the last three rows are initialized by EVA-CLIP.
-Method PC-459 A-150 PC-59 Stuff
-OpenAI 9.3 17.6 27.4 18.0
-CAT-Seg 19.6 34.1 57.3 45.6
-MaskCLIP++ 21.3 34.8 59.3 47.2
-EVA02 11.1 20.1 26.0 18.3
-CLIPSelf 13.0 22.9 39.1 27.1
-MaskCLIP++ 24.2 37.8 62.3 50.6
-Open-vocabulary panoptic and instance segmentation.
-Following the settings of previous works, we fine-tune CLIP on COCO Panoptic (Lin et al., 2014).
-All methods are evaluated on ADE20K, with PQ measuring panoptic segmentation performance and AP measuring instance segmentation performance.
-We use the same mask generator as FC-CLIP (Yu et al., 2024) and do not use the ensemble strategy ($\gamma=0.0$).
-Table 6 shows that our method achieves the best performance.
-Inference with the same masks.
-Considering that some competitive methods in Table 5 do not explicitly use a mask generator and fine-tune CLIP in alternative ways, we test the mask classification performance of CLIPs from various sources under the same generated masks.
-Since the PSM module we trained may not be compatible with their CLIP, we remove the PSM during inference.
-The ensemble strategy with the mask generator is also excluded.
-As shown in Table 7, our fine-tuned CLIPs outperform ones from other methods in mask classification.
-Inferece without mask generators.
-Because the goal of our fine-tuning is to enhance the mask classification ability of CLIP, which indirectly enhances the local representation between different patch representations of the CLIP.
-As shown in Table 8, after applying to some methods that do not require training and additional mask generators, further performance improvements can be obtained.
-Table 8: Performance improvements achieved when combined with several unsupervised open-vocabulary semantic segmentation methods that do not rely on a mask generator.
-OpenAI CLIP ViT-B/16 are used.
-“Citys” means cityscapes dataset (Cordts et al., 2016).
-Method A-847 PC-459 A-150 PC-59 Stuff Citys
-SCLIP (Wang et al., 2024a) - 7.7 16.1 34.2 22.4 32.2
-SCLIP w/ MaskCLIP++ - 9.5 19.3 40.8 29.0 35.9
-ClearCLIP (Lan et al., 2024a) - 7.9 16.7 36.0 23.9 30.0
-ClearCLIP w/ MaskCLIP++ - 9.6 19.2 40.8 30.4 33.5
-ProxyCLIP  (Lan et al., 2024b) 7.4 8.4 20.2 39.1 26.5 38.1
-ProxyCLIP w/ MaskCLIP++ 8.0 9.9 23.2 44.6 34.2 42.4
-Trade-off between training cost and performance.
-Figure 4 shows the trade-off between training cost and performance of different open-vocabulary segmentation methods.
-As our method does not use the mask generator or teacher model during training but only a small number of GT masks instead of a large number of generated masks and has only one classification loss, it converges quickly, has lower memory usage, and achieves better performance.
-Data efficiency.
-Since the goal of our fine-tuning is to transfer CLIP’s open-vocabulary recognition capabilities to mask regions, rather than learning category-specific pixel groupings from segmentation datasets, our method can effectively fine-tune using a smaller amount of image segmentation data.
-We randomly sampled varying proportions of data from the COCO-Stuff training set for fine-tuning, using the same configuration as the base model in Table 5.
-As shown in Table 9, even with only 0.1% of the data (approximately 118 images with annotations), the fine-tuned model still demonstrates competitive performance.
-Table 9: The impact of the amount of data used for fine-tuning.
-“data size” indicates the proportion of data used for fine-tuning relative to the original COCO-Stuff training set, and the fine-tuned model is CLIP-ConvNeXt-B.
-Data size A-847 PC-459 A-150 PC-59 Stuff
-100% 14.5 18.7 35.4 59.1 47.1
-10% 14.2 19.0 34.9 59.2 47.2
-5% 13.9 18.7 35.2 59.6 47.3
-1% 13.8 19.1 35.1 58.5 47.0
-0.1% 12.8 18.6 34.1 58.2 46.1
-Figure 4: Trade-off between training cost and performance of different OVS methods. Radius size represents the memory usage during training. Using CLIP with base size for comparison.
-Qualitative comparison of segmentation results.
-Figure 7 shows the qualitative results on the Open-Vocabulary Semantic Segmentation task.
-Using the mask generator from MAFT+ (Jiao et al., 2024), we compare the segmentation results of the original CLIP, MAFT+, and our method.
-Compared to the original CLIP, our method achieves more accurate recognition of local images.
-Compared to MAFT+, our method retains more of the original CLIP’s knowledge without generating too many erroneous biases.
-Our method also predicts relatively correct semantics for some objects that are not annotated in the GT.
-5 Conclusions
-In this paper, we point that using generated masks hinders CLIP’s regional image-text alignment in previous OVS work.
-And there is limited room for improving mask generation compared to mask classification.
-We also reveal that attempts to update CLIP’s embedding vectors disrupt its original pre-alignment properties, resulting in new alignments that function only on the training data.
-Based on these analysis, we propose the MaskCLIP++, which utilizes ground truth masks as prior to fine-tune CLIP, and models similarity under a consistency alignment constraint to mitigate overfitting.
-Our method significantly enhances CLIP’s mask recognition capabilities across various architectures.
-By integrating mask generators from prior work, we achieve comparable or superior performance on semantic segmentation, panoptic segmentation, and instance segmentation tasks with lower training costs.
-Overall, we adapt CLIP more effectively to open-vocabulary segmentation tasks through the use of image segmentation data and lightweight fine-tuning.
-Appendix
-Appendix A Limitations and Future work
-First, the image masks must be downsampled to match the size of CLIP’s final feature map, which undergoes a downsampling factor of 14 to 32 times.
-Consequently, some generated small object masks may not be recognizable by MaskCLIP++.
-Second, the mask generators employed during inference incur negligible computational overhead.
-As shown in Table 3, heavier backbones seem to enhance classification capabilities of their own ($P_{s}$), but do not significantly improve open-vocabulary segmentation performance when combined with our fine-tuned CLIP ($P_{c}$).
-Therefore, lightweight, category-agnostic mask generators continue to represent a promising research direction.
-Last, our reliance on GT masks somewhat limits our ability to scale up at the data level. Finding ways to utilize more image-text pairs or image-mask pairs for fine-tuning remains an open question.
-Appendix B Datasets of training and evaluation
-The image segmentation datasets we used to train are COCO Panoptic (Lin et al., 2014) or COCO-Stuff (Caesar et al., 2018).
-For evaluation, we utilize the ADE20K (Zhou et al., 2017), PASCAL VOC (Everingham et al., 2010), and PASCAL Context (Mottaghi et al., 2014) datasets.
-A brief introduction to these datasets is provided below.
-COCO Panoptic.
-It contains 164K images with 133 annotated classes, comprising 80 “thing” categories and 53 “stuff” categories.
-The “thing” categories are annotated at the instance level, with different instances of the same category using separate masks within a single image.
-In contrast, the “stuff” categories are semantically annotated, where all regions belonging to the same category share a single mask in an image.
-COCO Stuff. It shares the same training images with COCO Panoptic. These images are annotated for semantic segmentation into 171 categories.
-ADE20K.
-It contains 2,000 images for validation with two types of annotations: one for panoptic or semantic segmentation with 150 categories (dubbed as A-150), and the other for semantic segmentation with 847 categories (dubbed as A-847).
-Among the 150 categories defined in ADE20K, 64 appear in COCO Panoptic, while 86 do not.
-The names of these categories are detailed in Table 10.
-PASCAL VOC. It includes 1,449 images for testing with 20 annotated foreground classes for semnantic segmentation.
-PASCAL Context.
-It is a dataset for semantic understanding that contains 5,000 validation images.
-Two versions are employed for semantic segmentation: one with 59 frequently used classes (referred to as PC-59) and another with the complete set of 459 classes (referred to as PC-459).
-Table 10: Categories in ADE20K that are present in COCO Panoptic (seen) or are absent from COCO Panoptic (unseen).
+We set $\gamma=0.1$ for our model when integrating with the mask generator.
+However, as demonstrated in Tab.
+15, our’s performance enhancement stems primarily from improved mask classification capability rather than this integration strategy.
+Appendix B Dataset Details
+Dataset #Classes #Images
+Train
+COCO Stuff [5] 171 118280
+COCO Panoptic [36] 133 118287
+Eval
+COCO Stuff 171 5000
+COCO Panoptic 133 5000
+ADE20K [80] 150 / 847 2000
+PASCAL Context [43] 59 / 459 5000
+PASCAL VOC [17] 20 (w/o background) 1449
+Cityscapes [13] 19 500
+Eval (MESS - General)
+BDD100K [75] 19 1000
+Dark Zurich [49] 20 50
+MHP v1 [32] 19 980
+FoodSeg103 [66] 104 2135
+ATLANTIS [16] 56 1295
+DRAM [12] 12 718
+Eval (MESS - Earth)
+iSAID [64] 16 4055
+ISPRS Potsdam [56] 6 504
+WorldFloods [24] 3 160
+FloodNet [48] 10 5571
+UAVid [41] 8 840
+Eval (MESS - Medical)
+Kvasir-Inst. [45] 2 118
+CHASE DB1 [18] 2 20
+CryoNuSeg [42] 2 30
+PAXRay-4 [50] 2 180
+Eval (MESS - Engineer)
+Corrosion CS [3] 4 44
+DeepCrack [37] 2 237
+ZeroWaste-f [2] 5 929
+PST900 [53] 5 288
+Eval (MESS - Agriculture)
+SUIM [23] 8 110
+CUB-200 [58] 201 5794
+CWFID [21] 3 21
+Table 10: Statistics of datasets used in our experiments.
+Tab.
+10 provides detailed statistics for the datasets used in our experiments.
+Tab.
+11 lists the categories in ADE20K that are present in COCO Panoptic (seen) or absent from COCO Panoptic (unseen).
+This split is used in our experiments to evaluate the generalization capability in Sec.
+4.3.
 Split Name
@@
 canopy, flag, bar, ball
+Table 11: Categories in ADE20K that are present in COCO Panoptic (seen) or are absent from COCO Panoptic (unseen).
+Method BDD100K Dark Zurich MHPv1 FoodSeg103 ATLANTIS DRAM
+OpenAI CLIP 44.6 27.5 48.6 57.1 58.1 89.3
+EVA02 CLIP 44.7 34.3 49.2 61.4 61.0 91.0
+CLIPSelf 51.8 30.6 61.0 47.4 59.1 65.3
+CAT-Seg 59.3 40.3 49.8 54.6 65.3 94.6
+MaskCLIP++ 67.2 48.3 64.9 64.7 64.5 98.0
+(a) General
+Method iSAID ISPRS Potsdam WorldFloods FloodNet UAVid
+OpenAI CLIP 59.2 65.5 65.9 65.9 48.9
+EVA02 CLIP 61.0 68.4 71.4 59.6 42.2
+CLIPSelf 43.6 41.7 42.4 38.3 62.6
+CAT-Seg 59.8 76.9 71.0 62.3 64.4
+MaskCLIP++ 60.5 80.0 75.1 63.0 64.4
+(b) Earth Monitoring
+SUIM CUB-200 CWFID
+55.5 54.7 42.8
+57.6 60.0 39.7
+56.0 51.1 61.9
+72.9 47.2 73.0
+73.2 40.8 82.5
+(c) Agriculture and Biology
+Method Kvasir Inst. CHASE DB1 CryoNuSeg PAXRay-4
+OpenAI CLIP 93.2 50.0 50.0 50.8
+EVA02 CLIP 95.0 50.0 50.0 51.7
+CLIPSelf 72.0 50.0 50.0 50.0
+CAT-Seg 89.8 50.0 50.0 58.1
+MaskCLIP++ 99.2 50.0 50.0 99.4
+(d) Medical Sciences
+Corrosion CS DeepCrack ZeroWaste-f PST900
+25.5 69.2 32.5 19.4
+25.0 58.9 33.0 18.4
+31.6 76.8 45.2 27.5
+21.6 61.0 30.2 26.7
+24.4 85.7 57.4 33.8
+(e) Engineering
+Table 12: Mask accuracy on MESS datasets [4].
 Appendix C Additional Experiments
-C.1 Additional Illustrations
-Implementation details of Figure 1 We use mask generation and mask classification as the inference modes, with Figure 1(a) and Figure 1(b) presenting the performance upper bounds of the ideal mask generator and ideal mask classifier, respectively.
-The ideal mask generator is implemented by directly using the ground truth masks.
-To implement the ideal mask classifier, we compute the maximum matching between the generated masks and ground truth masks based on IoU, then assign the corresponding GT categories to the matched masks and combine them into the output segmentation maps.
-Figure 5: A toy example: Because the inconsistency alignment between $s$ (similarity before $\Theta$) and $r$ (similarity after $\Theta$), after a round of gradient descent, $s$ is further from the target $\hat{s}$.
-Phenomenon of inconsistency alignment.
-We first use a toy example to illustrate the potential consequences of destroying alignment consistency.
-Assume that we use the parameter $\Theta$ to update the mask embeddings, changing the original similarity from $s$ to $r$, and then optimize towards the target.
-However, due to the lack of explicit constraints on $\Theta$, there may be a situation as shown in Figure 5, where the order of $s$ and $r$ is inconsistent, resulting in correct optimization on $r$ but incorrect optimization on $s$.
-For simplicity, $\mathcal{L}$ in Figure 5 denotes an L1 loss without softmax, and only mask embeddings $m$ are updated in one round of gradient descent.
-Futuremore, we use Table 12 as a supplement to Table 2 to illustrate the extent to which different PSM structures destroy the consistency alignment constraint.
-For models fine-tuned with various PSMs, we test by retaining or removing the PSM module, and compare the performance differences to assess whether the alignment consistency is maintained before and after the PSM intervention.
-The results show that only the three blue PSM structures largely satisfy the consistency alignment constraint, while the three red PSM structures severely disrupt it.
-Table 11: The phenomenon of consistency alignment.
-PSM test mIoU(S) mIoU(U) mIoU
-- - 32.7 16.3 23.3
-$\langle\operatorname{Linear}(E_{m}),E_{t}\rangle$ 37.7 20.2 27.7
-32.4 1.9 14.9
-$\langle E_{m},\operatorname{Linear}(E_{t})\rangle$ 37.6 20.3 27.6
-32.5 1.8 14.9
-$\operatorname{Linear}\langle E_{m},E_{t}\rangle$ 45.2 24.6 33.4
-45.4 25.7 34.1
-$\langle E_{m},\operatorname{Norm}(P_{t})\rangle$ 39.5 21.4 29.1
-31.9 4.5 16.2
-$\langle E_{m},\operatorname{Norm}(E_{t})\rangle$ 45.8 24.4 33.5
-46.2 25.2 34.1
-$\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle$ 44.0 24.1 32.6
-46.3 26.7 35.1
-Table 12: Performance of MaskCLIP++ on closed-vocabulary image segmentation tasks.
-CLIP arch. Dataset mIoU PQ AP
-ConvNeXt-B COCO-Stuff 47.1 - -
-ViT-L/14 COCO-Stuff 51.0 - -
-ConvNeXt-B COCO-Panoptic 57.6 46.1 34.4
-ViT-L/14 COCO-Panoptic 63.3 52.3 42.3
-The performance of closed-vocabulary image segmentation.
-We present the performance of MaskCLIP++ on closed-vocabulary image segmentation in Table 12, with the same configuration as in Table 5 and Table 6, respectively.
-C.2 Additional Ablations
-(a) ConvNeXt-B
-(b) ConvNeXt-L
-(c) ViT-B/16
-(d) ViT-L/14
-Figure 6: Under various input sizes of different CLIP models, our method (blue bars) has significant improvements over the original CLIP (red bars).
-The input size of the mask generator remains unchanged.
-All models are fine-tuned on COCO Panoptic and evaluated on ADE20K with mIoU reported.
-Impact of CLIP input size.
-Considering that input image size often has a significant impact on the mIoU of semantic segmentation models, we compared the performance of our method with the original CLIP under different input sizes.
-As shown in Figure 6, with a fixed input size for the mask generator, our fine-tuned CLIP consistently outperforms the original CLIP across various input sizes for open-vocabulary semantic segmentation.
-Overall, segmentation performance initially increases and then decreases as input size grows.
-Since our fine-tuned CLIP can better recognize local image features, the optimal input size tends to be smaller than that of the original CLIP.
-Our framework allows us to maintain the input size for the mask generator while reducing the input size for mask recognition, thereby enhancing performance and reducing computational costs.
-Impact of integrating image context information into mask representation.
-Averaging CLIP features over perfectly precise mask regions is not conducive to extracting mask representations with stronger recognition capabilities, as recognition tasks typically require contextual information and focused attention on target internal regions.
-Here, we distinguish how we address this common issue using CLIP’s original pooling method.
-For CLIP models utilizing Attention Pooling (e.g., ViT-B/16), the hyperparameter $\alpha$ controls the degree of contextual information incorporated in the mask representation.
-As shown in Table 1313(b), smaller $\alpha$ values, indicating more extensive use of contextual information, lead to better performance.
-For CLIP models employing global average pooling (e.g.
-ConvNeXt-B), as shown in Table 1313(a), using global image embeddings $E_{i}$ or patch embeddings $E_{p}$ to construct pseudo-text embeddings $P_{t}$ in the PSM stage enhances the utilization of image context information to some extent, thereby improving performance.
-Choice of Loss.
-Table 15 presents the performance differences observed when fine-tuning CLIP ConvNeXt-B with various loss functions.
-Training with COCO Panoptic (Lin et al., 2014) and evaluating on ADE20K (Zhou et al., 2017).
-When the training labels for a single image include multiple instances of the same category, such as pixels or instance masks, we re-weight the loss by categories to prevent bias toward more common classes during optimization.
-We adopt a class-reweighted cross-entropy loss for all our experiments in the main text.
-Position of the mask prior insertion.
-Due to the repetitive nature of the ViT architecture, the position of mask prior insertion in CLIP models based on such architectures is not absolute.
-Following the approach of Xu et al.
-(2023c) for constructing attention masks, we can insert the mask prior at an earlier position than the last attention block.
-Table 15 takes ViT-B/16 as an example, demonstrating that inserting the mask prior at the third-to-last layer yields better performance.
-Therefore, in Table 4, the mask prior is inserted at the third-to-last layer for all ViT-based CLIP models.
-Table 13: Impact of integrating image context information into mask embeddings $E_{m}$. The implementation varies based on the original CLIP’s pooling method.
-(a) The case of CLIP using global average pooling, taking ConvNeXt-B as an example.
-$\Psi$ is a placeholder for image context information, and $P_{t}$ defined in equation 3 can be rewritten as $P_{t}=\operatorname{Attn}(q(E_{t}),k(\Psi),v(\Psi))$.
-$P_{t}$ interacts with $E_{m}$ in the PSM stage to increase the utilization of image context information.
-Method in PSM Image Context mIoU
-No $\Phi$ 34.1
-$\Phi=E_{i}$ 34.5
-$\Phi=E_{p}$ 35.1
-(b) The case of CLIP using attention pooling, taking ViT-B/16 as an example.
-$\alpha$ is defined in equation 2.
-The bigger $\alpha$ is, the less the semantic similarity of the image context is utilized.
-The learnable $\alpha$ is initialized from $e^{-5}$ and learned in the logarithmic space.
-Method in CLIP Image Context mIoU
-$\alpha=e^{5}$ 31.1
-$\alpha=e^{-5}$ 33.5
-learnable $\alpha$ 33.8
-Table 14: Choice of Loss. “Re.” means the loss is reweighted by categories. CLIP ConvNeXt-B is used.
-Prior Loss Re. mIoU(S) mIoU(U) mIoU
-Pixel CE 39.5 21.2 29.0
-CE 44.3 23.4 32.8
-Instance Mask BCE 45.7 24.5 33.5
-BCE 46.3 25.3 34.3
-CE 45.5 26.0 34.4
-CE 46.3 26.7 35.1
-Table 15: The impact of the position of mask prior insertion for ViT-B/16 CLIP. “Prior pos.” denotes the insert position relative to the total number of attention blocks.
-Prior pos. mIoU PQ
--1 33.3 24.1
--2 33.5 24.5
--3 33.8 24.4
--4 33.5 24.4
-Qualitative results.
-Figure 7 compares the visual results of open-vocabulary semantic segmentation using different methods with the same mask generator.
-We show more qualitative results of semantic- and instance level of open-vocabulary segmentation in Figure 8.
-Image                    CLIP                MAFT+              MaskCLIP++ GT
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Figure 7: Visualizations of open-vocabulary semantic segmentation on ADE20K (Zhou et al., 2017).
-From left to right are the input image, the result using the original CLIP, the result of MAFT+ (Jiao et al., 2024), the result using MaskCLIP++, and the ground truth.
-The same mask generator is used for CLIP, MAFT+ and MaskCLIP++.
+C.1 Mask accuracy on MESS
+Tab.
+12 complements Tab.
+1 by listing the mask accuracy of different models on each dataset in MESS [4].
+MaskCLIP++ achieves better maskAcc on 17 out of 22 datasets, providing compelling empirical evidence that the proposed method significantly enhances mask classification accuracy while simultaneously preserving the inherent generalization capabilities of the original CLIP.
+However, note that in specialized fine-grained classification tasks like the bird-oriented CUB-200 dataset [58], where category labels are semantically close and images usually have a single foreground subject, fine-grained category domain-specific knowledge impacts performance more than localized perception.
+Thus, CLIP models fine-tuned on COCO generally perform worse than the original CLIP in such cases.
+C.2 Effectiveness of mask tuning on different CLIP architectures
+To validate the generalizability of our approach across varying CLIP architectures, we conduct fine-tuning experiments on the COCO Panoptic dataset [36] using distinct CLIP architectures and report the mask classification accuracy of our method as documented in Tab.
+13.
+Arch. A-847 PC-459 A-150 PC-59 Stuff
+RN50x16 24.1 37.5 50.4 73.3 54.1
+ConvNeXt-B 30.0 42.1 56.4 76.3 55.6
+ConvNeXt-L 33.0 50.8 61.0 78.7 59.0
+ConvNeXt-XXL 38.7 56.1 64.6 81.5 62.5
+ViT-B/16 29.3 45.7 58.3 78.6 58.0
+ViT-L/14 37.5 56.9 66.0 83.2 63.6
+Table 13: Mask accuracy of MaskCLIP++ on different architectures fine-tuned on COCO Panoptic [36].
+C.3 Other ablation studies
+Effect of Fuser depth.
+The positioning of mask insertions exerts non-negligible effects on model performance.
+As delineated in Sec.
+3.2, the quantity of Transformer blocks in the Fuser module can be mathematically formulated as $L-K$.
+Tab.
+14(a) presents a systematic analysis of how varying Fuser depths impact OVS performance.
+Considering the computational overhead that scales linearly with mask quantity when increasing Fuser depth, we empirically select $L-K=2$ as the default configuration.
+$\bm{L-K}$ mIoU PQ
+1 33.3 24.1
+2 33.5 24.5
+3 33.8 24.4
+4 33.5 24.4
+(a) Effect of Fuser depth.
+$\bm{\alpha}$ mIoU
+$e^{5}$ 33.1
+$e^{-5}$ 33.5
+learnable 33.8
+(b) Effect of $\alpha$.
+Table 14: Ablation studies on the effect of Fuser depth and $\alpha$.
+Effect of $\alpha$.
+As defined in Eq.
+2, the bigger $\alpha$ is, the less the semantic similarity of the image context is utilized.
+As $\alpha\rightarrow+\infty$, $E_{m}^{(l+1)}$ degenerates to the mask average pooling result of $F^{(l)}$, thereby inadequately leveraging contextual information beyond the masked regions.
+Conversely, when $\alpha=0$, $E_{m}^{(l+1)}$ reduces to the CLS token after $l$-th layer.
+As shown in Tab.
+14(b), selecting a smaller $\alpha$ value yields superior performance.
+Therefore, we initialize $\alpha$ as $e^{-5}$ and adaptively adjust it in the logarithmic space during training.
+Effect of classification ensemble.
+To demonstrate that the performance gains of our method primarily stem from enhanced mask classification capabilities, we present in Tab.
+15 the comparative analysis of implementing versus omitting the ensemble strategy.
+Method Ensemble A-847 PC-459 A-150 PC-59
+FC-CLIP 12.1 12.7 27.4 42.8
+14.8 18.2 34.1 58.4
+MaskCLIP++ 15.3 21.4 36.8 62.4
+15.4 21.3 37.1 62.6
+Table 15: Effect of classification ensemble.
 Image                          Semantic Segmentation                Panoptic Segmentation Instance Segmentation
@@
 Refer to captionRefer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to captionRefer to caption
 Figure 7: More visualizations of semantic- and instance-level of open-vocabulary segmentation.
-We use mask generator from FC-CLIP (Yu et al., 2024) with ConvNeXt-L backbone.
-And use CLIP ViT-L/14 fine-tuned on COCO Panoptic (Lin et al., 2014).
-Images are from ADE20K (Zhou et al., 2017).
+We use mask generator from FC-CLIP [76] with ConvNeXt-L backbone.
+And use CLIP ViT-L/14 fine-tuned on COCO Panoptic [36].
+Images are from ADE20K [80].
 The confidience threshold for instance segmentation is 0.3.